### v2

# Factorizable joint shift revisited

Abstract
Factorizable joint shift (FJS) represents a type of distribution shift (or dataset shift) that comprises both covariate and label shift. Recently, it has been observed that FJS actually arises from consecutive label and covariate (or vice versa) shifts. Research into FJS so far has been confined to the case of categorical labels. We propose a framework for analysing distribution shift in the case of a general label space, thus covering both classification and regression models. Based on the framework, we generalise existing results on FJS to general label spaces and present and analyse a related extension of the expectation maximisation (EM) algorithm for class prior probabilities. We also take a fresh look at generalized label shift (GLS) in the case of a general label space.
Keywords: Distribution shift, dataset shift, factorizable joint shift, generalized label shift, domain adaptation, EM algorithm.
1 Introduction
Classification and regression models may work very well on their respective training datasets but fail utterly when being deployed in test datasets. Such failure can be caused by distribution shift (also known as dataset shift) between the training and test datasets. For this reason, distribution shift and domain adaptation (a notion comprising techniques for tackling distribution shift) has been a major research topic in machine learning for some time. This paper takes the perspective of Kouw and Loog (2021) and studies the case where feature observations from the test dataset are available for analysis but observations of labels are missing.
Under these circumstances, without any assumptions on the nature of the distribution shift between the training and test datasets meaningful prediction of the labels in the test dataset or of their distribution is not feasible. See Kouw and Loog (2021) for a survey of approaches to domain adaptation and their related assumptions. Arguably, covariate shift (also known as population drift) and label shift (also known as prior probability shift or target shift) are the most popular specific distribution shift assumptions, both for their intuiveness as well as their computational manageability. However, exclusive covariate and label shift assumptions have been criticised for being insufficient for common domain adaptation tasks (e.g. Tachet des Combes et al., 2020).
He et al. (2021) proposed the notion of factorizable joint shift (FJS) as an assumption for distribution shift that covers both covariate and label shift. Recently, Dong et al. (2025) observed that FJS actually may be described as combination of label and covariate shifts, thus making the assumption more plausible. Tasche (2022) analysed FJS for categorical label spaces with regard to the potential and limitations of domain adaptation based on this assumption. This paper updates the analyses of Tasche (2022) for FJS in the case of a general label space. More precisely, the main contributions of this paper to the subject of distribution shift are the following:
  • Pointing out the building blocks of non-specific distribution shift with general label spaces (including Euclidean, categorical and mixed Euclidean-categorical label spaces) in order to facilitate the demarcation and analysis of specific types of distribution shift.
  • Generalising the characterisation of FJS by Tasche (2022) such that both classification and regression problems are covered.
  • Presenting a generalisation to general label spaces of the expectation maximisation (EM) algorithm for estimating target prior class probabilities by Saerens et al. (2002) and showing that it is fit for purpose.
  • Generalising a result by He et al. (2021) on the relation between generalized label shift (GLS) and FJS.
The setting and notation for this paper are set out in the first two sub-sections of Section 2, followed by a framework of general assumptions and results on non-specific distribution shift in the remaining sub-sections of Section 2. Section 3 comprises analyses and results regarding FJS as well as approaches to estimating the characterics of FJS, including the presentation and analysis of a general EM algorithm in Sub-section 3.2.3. The relation between GLS and FJS is analysed in Section 4. Section 5 concludes the paper, with a summary and suggestions for further research. Appendix A presents a description of the classification approach to density estimation while Appendix B contains the proofs for the results in Sub-section 3.2.3.
2 The building blocks of distribution shift
We study distribution shift related to classification and regression problems in a probabilistic setting with feature (also known as input, explanatory or independent) variables $X$ and label (also known as output, response or dependent) variables $Y$. We assume that a joint source (also known as training) probability distribution $P=P_{(X,Y)}$ of $X$ and $Y$ already has been learnt. Now we are presented with a joint target (also known as test) distribution $Q=Q_{(X,Y)}$ of $X$ and $Y$ which is possibly different to $P$ and only partially observable. Lack of observations from the target distribution $Q$ and more generally lack of information about $Q$ may become an issue in the presence of distribution shift (Zhang et al., 2024, Section 4.9). The problem then is to infer information about the posterior (also known as conditional) distribution $Q_{Y|X}$ of $Y$ given $X$ under the target distribution $Q$ from the source distribution $P$ despite a lack of joint observations of $(X,Y)$ under $Q$. In Section 3 below, we analyse the two cases where
  • (A)
    both of the target marginal distributions $Q_{X}$ and $Q_{Y}$ of the feature and label variables respectively are known but neither of the posterior distributions $Q_{X|Y}$ and $Q_{Y|X}$, and
  • (B)
    the target marginal distribution $Q_{X}$ of the feature variable is known but neither the marginal distribution $Q_{Y}$ of the label variable nor any of the posterior distributions $Q_{X|Y}$ and $Q_{Y|X}$.
The problem phrased in (B) is an unsupervised domain adaptation problem where a classification or regression function is learnt on a completely known dataset in a supervised manner and is then adapted to another dataset without label information in an unsupervised manner (Kouw and Loog, 2021, Section 6.7). Problem (A) where only joint observations of features and label are missing has less frequently been looked at in the machine learning literature. It is akin to copula estimation (Durante and Sempi, 2016) and as such has some practical relevance (see e.g. Jaworski et al., 2013; Tasche, 2025). Problem (A) is investigated in this paper mainly because its solutions can inform approaches to solutions for Problem (B).
In preparation for the discussion of FJS in Section 3, in this section we assume that the source distribution $P$ and the target distribution $Q$ are both completely known. This assumption is reasonable because
  • it helps identifying building blocks of the distributions which can be estimated from observations under appropriate assumptions,
  • it helps identifying building blocks of the distributions for which reasonable assumptions can be made (based on exogenous information, on theoretical considerations etc.), and
  • knowledge of the building blocks can inform generative approaches to distribution shift when datasets need to be generated by simulation.
2.1 Setting
The setting for this paper is a standard setting for domain adaptation and therefore, with the exception of some details of the notation, similar to the setting described for example in the first paragraph of Section 2 of Azizzadenesheli (2022).
  • Each instance (also known as object) $\omega$ is an element of a set $\Omega$ which is called instance space. Each instance $\omega$ has features which are represented collectively as $x\in\Omega_{X}$ and a label $y\in\Omega_{Y}$. $\Omega_{X}$ and $\Omega_{Y}$ are called feature space and label space respectively. We are interested in predicting [^1] If $\Omega_{Y}$ is a finite or countably infinite set the prediction task is called classification. Otherwise it is called regression. $y$ from $x$ and assume that any information available for the prediction is captured by $x$. Therefore, instances $\omega$ may be identified with combinations of $x$ and $y$ such that $\omega=(x,y)$. Consequently, without loss of generality it can be assumed that $\Omega=\Omega_{X}\times\Omega_{Y}$.
  • We call the projection $X:\Omega\to\Omega_{X},\omega\mapsto x=X(\omega)$ feature variable and the projection $Y:\Omega\to\Omega_{Y},\omega\mapsto y=Y(\omega)$ label variable.
  • Instances occur with different frequencies, for example instances represented by a combination $\omega_{0}=(x_{0},y_{0})$ of features and label might be observed more often than instances with a combination $\omega_{1}=(x_{1},y_{1})$. To reflect such different occurence frequencies and to together cover classification and regression as well as to properly deal with mixed features consisting of both continuous and categorical components, the setting for this paper is probabilistic and formulated in terms of measure theory (cf. Chapter 1 of Klenke, 2013).
  • The feature space $\Omega_{X}$ and the label space $\Omega_{Y}$ are parts of measurable spaces $(\Omega_{X},\mathcal{H})$ and $(\Omega_{Y},\mathcal{G})$ respectively where the $\sigma$-algebra (also known as information set, Holzmann and Eulert, 2014) $\mathcal{H}$ is a collection of subsets of $\Omega_{X}$ and the $\sigma$-algebra $\mathcal{G}$ is a collection of subsets of $\Omega_{Y}$. The related $\sigma$-algebra $\mathcal{F}$ of subsets of the instance space $\Omega$ then is defined as the $\sigma$-algebra generated by the feature variable $X$ and the label variable $Y$, i.e.
    $\mathcal{F}\ =\ \sigma(X,Y)\ =\ \sigma\bigl(X^{-1}(\mathcal{H})\cup Y^{-1}(\mathcal{G})\bigr)\ =\ \mathcal{H}\otimes\mathcal{G},$ (2.1)
    with $X^{-1}(\mathcal{H})=\{X^{-1}(H):H\in\mathcal{H}\}$ and $Y^{-1}(\mathcal{G})=\{Y^{-1}(G):G\in\mathcal{G}\}$. By definition of $\mathcal{F}$, $X$ is $\mathcal{H}$-$\mathcal{F}$-measurable and $Y$ is $\mathcal{G}$-$\mathcal{F}$-measurable.
  • The source distribution $P=P_{(X,Y)}$ is a joint distribution of the feature variable $X$ and the label variable $Y$ and as such defined on $(\Omega,\mathcal{F})$. In the following, the source distribution $P$ is assumed to be completely known. The source feature distribution $P_{X}$ is the image measure (or push-forward measure) of $P$ under $X$ and as such defined by $P_{X}[H]=P[X\in H]$, $H\in\mathcal{H}$. The source label distribution $P_{Y}$ is the image measure of $P$ under $Y$ and as such defined by $P_{Y}[G]=P[X\in G]$, $G\in\mathcal{G}$. Both $P_{X}$ and $P_{Y}$ are unconditional distributions of $X$ and $Y$ respectively in the sense that they do not take into account already observed realisations of the other variable. In the machine learning literature $P_{X}$ and $P_{Y}$ also are referred to as prior distributions.
  • The target distribution $Q=Q_{(X,Y)}$ is another joint distribution of the feature variable $X$ and the label variable $Y$ and as such also defined on $(\Omega,\mathcal{F})$. As mentioned above, in contrast to the source distribution, the target distribution $Q$ is assumed to be only partially known to an extent to be specified below in Section 3. The target feature distribution $Q_{X}$ is the image measure of $Q$ under $X$ and as such defined by $Q_{X}[H]=Q[X\in H]$, $H\in\mathcal{H}$. The target label distribution $Q_{Y}$ is the image measure of $Q$ under $Y$ and as such defined by $Q_{Y}[G]=Q[X\in G]$, $G\in\mathcal{G}$. $Q_{X}$ and $Q_{Y}$ also are called prior distribution.
  • The source distribution $P$ and the target distribution $Q$ are different, i.e. $P\neq Q$. In other words, there is a distribution shift (see e.g. Zhang et al., 2024) between source and target. Occasionally, non-specific distribution shift also is called dataset shift in the literature (Quiñonero-Candela et al., 2009; Moreno-Torres et al., 2012).
2.2 Notation
The following notation and technical terms are broadly the same as the notation and terms used in textbooks on probability theory like Bauer (1981) or Klenke (2013). The notation for conditional expected values, probabilities and distributions is intended to pin down a rigorous understanding of these concepts because sometimes they are only superficially or not at all explained in the machine learning literature (e.g. Hastie et al., 2009; Murphy, 2012).
Expected values.
For $\mathcal{F}$-Borel-measurable random variables $T:\Omega\to\mathbb{R},\omega\mapsto T(\omega)$, the expected value of $T$ with respect to $P$ is
$E_{P}[T]\ =\ \int_{\Omega}T(\omega)\,P(d\omega)\ =\ \int T\,dP$
if $\int\lvert T\rvert\,dP<\infty$ or $T$ is non-negative. In the same way, expected values are defined with respect to other probability spaces like $(\Omega_{X},\mathcal{H},P_{X})$, $(\Omega,\mathcal{F},Q)$ etc.
For example, for $\mathcal{H}$-Borel-measurable $T:\Omega_{X}\to\mathbb{R},x\mapsto T(x)$, the expected value of $T$ with respect to $P_{X}$ is
$E_{P_{X}}[T]\ =\ \int_{\Omega_{X}}T(x)\,P_{X}(dx)\ =\ \int T\,dP_{X}\ =\ E_{P}[T(X)]$
if $\int\lvert T\rvert\,dP_{X}<\infty$ or $T$ is non-negative.
Indicator function.
For a set $M$, the indicator function $\mathbf{1}_{M}$ is defined as $\mathbf{1}_{M}(m)=1$ if $m\in M$ and $\mathbf{1}_{M}(m)=0$ if $m\notin M$.
‘Almost surely’.
We use the shorthand $\mathbb{P}$-a.s.’ for $\mathbb{P}$-almost surely’ which is a shorter version of the phrase ‘with probability $1$ under the distribution $\mathbb{P}$’. Any of the probability measures encountered in the following can be substituted for $\mathbb{P}$.
Expected values conditional on $\sigma$-algebras.
For $\mathcal{F}$-Borel-measurable $T:\Omega\to\mathbb{R},\omega\mapsto T(\omega)$ and any sub-$\sigma$-algebra $\mathcal{C}\subset\mathcal{F}$, the expected value of $T$ conditional on $\mathcal{C}$ under $P$ is the $\mathcal{C}$-Borel-measurable random variable $E_{P}[T\,|\,\mathcal{C}]$ on $(\Omega,\mathcal{F},P)$ which is determined $P$-a.s. by the property
$E_{P}\bigl[E_{P}[T\,|\,\mathcal{C}]\,\mathbf{1}_{C}\bigr]\ =\ E_{P}[T\,\mathbf{1}_{C}],\quad\text{for all}\ C\in\mathcal{C},$
if $\int\lvert T\rvert\,dP<\infty$ or $T$ is non-negative.
For $F\in\mathcal{F}$, the probability of $F$ conditional on $\mathcal{C}$ under $P$ is defined by $P[F\,|\,\mathcal{C}]=E_{P}[\mathbf{1}_{F}\,|\,\mathcal{C}]$. Expected values and probabilities conditional on $\sigma$-algebras under other probability measures like the target distribution $Q$ are defined analogously.
Expected values conditional on random variables.
Let $T:\Omega\to\mathbb{R},\omega\mapsto T(\omega)$ be $\mathcal{F}$-Borel-measurable and non-negative or satisfy $\int\lvert T\rvert\,dP<\infty$ and let $V$ be a measurable function on $(\Omega,\mathcal{F})$ with values in a measurable space $(\Omega_{V},\mathcal{V})$. Then the $\mathcal{V}$-Borel-measurable real-valued function $E_{P}[T\,|V=\cdot]$ is the expected value of $T$ conditional on $V$ under $P$ if it holds for all $M\in\mathcal{V}$ that
$E_{P}[T\,\mathbf{1}_{\{V\in M\}}]\ =\ \int_{M}E_{P}[T\,|V=v]P_{V}(dv),$ (2.2)
with $\{V\in M\}=V^{-1}(M)$ and the shorthand notation $E_{P}[T\,|V=v]=E_{P}[T\,|V=\cdot](v)$ for $v\in\Omega_{V}$. Note that $E_{P}[T\,|V=\cdot]$ is determined $P_{V}$-a.s. by property (2.2).
For $F\in\mathcal{F}$, the probability of $F$ conditional on $V$ under $P$ is defined by $P[F\,|\,V=\cdot]=E_{P}[\mathbf{1}_{F}\,|\,V=\cdot]$. Expected values and probabilities conditional on random variables under other probability measures like the target distribution $Q$ are defined analogously.
The definitions of expected values conditional on $\sigma$-algebras and on random variables respectively are related through the following identity:
$E_{P}[T\,|\,\sigma(V)](\omega)\ =\ E_{P}[T\,|V=V(\omega)]\quad P\text{-a.s.\ for}\ \omega\in\Omega,$ (2.3)
where $\sigma(V)=V^{-1}(\mathcal{V})$ is the sub-$\sigma$-algebra of $\mathcal{F}$ generated by $V$.
Expected values and probabilities conditional on $\sigma$-algebras or on random variables are well-defined for all combinations of probability spaces and integrable or non-negative random variables whose conditional expected values are meant to be determined (Bauer, 1981, Sections 10.1 and 10.2; Klenke, 2013, Sections 8.2 and 8.3).
Definition 2.1 (Regular conditional distribution).
In the setting of Section 2.1, a mapping $K:\Omega_{X}\times\mathcal{G}\to[0,\infty)$ is called (regular) conditional distribution of $Y$ given $X$ under $P$ if it has the following three properties:
  1. (i)
    For fixed $G\in\mathcal{G}$, the function $K(\cdot,G):\Omega_{X}\to[0,\infty),x\mapsto K(x,G)$ is $\mathcal{H}$-Borel-measurable.
  2. (ii)
    For fixed $x\in\Omega_{X}$, the function $K(x,\cdot):\mathcal{G}\to[0,\infty),G\mapsto K(x,G)$ is a probability measure on $(\Omega_{Y},\mathcal{G})$.
  3. (iii)
    For all $G\in\mathcal{G}$, it holds that $P[Y\in G\,|\,X=x]=K(x,G)$ $P_{X}$-a.s. for $x\in\Omega_{X}$.
In the following, we use the notation $P_{Y|X}$ for mappings $K$ satisfying (i) to (iii) and write $K(x,G)=P_{Y|X=x}[G]$ to make it clear that the conditional probability from (iii) is represented with a conditional distribution.
Regular conditional distributions $P_{X|Y}$ also are defined by (i) to (iii), with swapped roles of $X$ and $Y$. Conditional distributions under the target distribution $Q$ are defined by (i) to (iii), with $Q$ substituted for $P$.
In general, the existence of regular conditional distributions can be guaranteed only if certain conditions on the image of the conditioned variable – $(\Omega_{Y},\mathcal{G})$ in Definition 2.1 – are satisfied (Bauer, 1981, Section 10.3; Klenke, 2013, Section 8.3). If however the label variable $Y$ is real-valued, the regular conditional distribution $P_{Y|X}$ exists, without any restrictions on the feature variable $X$ (Klenke, 2013, Theorem 8.28).
In Section 2.4 below, we introduce Assumption 2.9 that suffices to guarantee the existence of both $P_{Y|X}$ and $P_{X|Y}$ and to some extent can be checked empirically. The need to have regular conditional distributions is context-dependent. In the following, it is mentioned explicitly when regular conditional distributions $P_{X|Y}$ or $P_{Y|X}$ in the sense of Definition 2.1 are supposed to exist.
Regular conditional distributions are required in particular if integration with respect to $P_{X|Y}$ or $P_{Y|X}$ is involved, as for Fubini’s theorem for conditional distributions (Klenke, 2013, Theorem 14.26):
$E_{P}[T(X,Y)]\ =\ \int\left(\int T(x,y)P_{X|Y=y}(dx)\right)P_{Y}(dy)\ =\ \int\left(\int T(x,y)P_{Y|X=x}(dy)\right)P_{X}(dx),$ (2.4a)
for all $\mathcal{F}$-Borel measurable $P$-integrable or non-negative functions $T=T(X,Y):\Omega\to\mathbb{R}$ if the related regular conditional distributions $P_{X|Y}$ and $P_{Y|X}$ respectively exist.
As a consequence of (2.4a) expected values conditional on the features $X$ or on the labels $Y$ are related to conditional distributions as follows
$\begin{split}E_{P}[T(X,Y)\,|\,Y=y]&\ =\ \int T(x,y)P_{X|Y=y}(dx),\ P_{Y}\text{-a.s.\ for all}\ y\in\Omega_{Y},\\ E_{P}[T(X,Y)\,|\,X=x]&\ =\ \int T(x,y)P_{Y|X=x}(dy),\ P_{X}\text{-a.s.\ for all}\ x\in\Omega_{X},\end{split}$ (2.4b)
under the same conditions as for (2.4a).
2.3 Linking source and target distributions in the face of distribution shift
Zhang et al. (2013) commented on the problem of how to deal with distribution shift: “If the data distribution changes arbitrarily, training data would be of no use to make predictions on the test domain. To perform domain adaptation successfully, relevant knowledge in the training (or source) domain should be transferred to the test (or target) domain.”
In this paper, knowledge transfer from source distribution $P$ to target distribution $Q$ is expressed through a density, as stated in the following main assumption of the paper.
Assumption 2.2 (Absolute continuity of target distribution).
In the setting of Section 2.1, $Q$ is absolutely continuous with respect to $P$, i.e.
[…]
$f=\frac{dQ}{dP}$ denotes the Radon-Nikodym derivative [^2] The existence of $f$ follows from the Radon-Nikodym theorem (Klenke, 2013, Corollary 7.34). of $Q$ with respect to $P$, i.e. $f$ is a density of $Q$ with respect to $P$ in the sense that $f$ is $\mathcal{F}$-Borel measurable and non-negative and it holds that
$Q[F]\ =\ E_{P}[f\,\mathbf{1}_{F}]\quad\text{for all}\ F\in\mathcal{F}.$ (2.5b)
In the literature the ‘transfer’ of information from $P$ to $Q$ by means of (2.5b) is sometimes called ‘importance reweighting’ and the density $f$ is referred to as ‘importance weights’ (e.g. Zhang et al., 2013; Kouw and Loog, 2021). More often than not, in the literature the reference measure for the density $f$ remains undisclosed or appears to be the Lebesgue measure (e.g. Zhang et al., 2013; Lipton et al., 2018). A notable exceptions are Scott (2019) who however proposes a reference measure $\mu$ that is available only in the case of a categorical label space, and Azizzadenesheli (2022).
Observe that condition (2.5a) is equivalent to stating that $Q[F]>0$ for $F\in\mathcal{F}$ implies $P[F]>0$. Hence condition (2.5a) may be interpreted as ‘the source distribution $P$ is richer with information than the target distribution $Q$’. Actually, this is a rather strong assumption such that there is a host of literature on approaches to domain adaptation without Assumption 2.2, including the seminal paper by Ben-David et al. (2006). See also Johansson et al. (2019) for analyses of the consequences of renouncing Assumption 2.2.
For the rest of the paper, we suppose that Assumption 2.2 is true. If there are independent joint feature-label observations from both the source and the target distributions, the density $f$ from the assumption can be estimated. Besides giving a short overview of the available estimation methods, Filipović and Schneider (2025) present methods based on reproducing kernel Hilbert spaces (RKHS) which are applicable under mild conditions. An approach based on probabilistic binary classification was proposed by Qin (1998). This approach is attractive because it allows for measuring the similarity of the two distributions in question with tools like the area under the curve (AUC) or the test statistic of the Kolmogorov-Smirnov test. For the readers’ convenience, a description of the classification approach is given in Appendix A below.
Proposition 2.3 records some immediate consequences of Assumption 2.2.
Proposition 2.3.
Let $T=T(X,Y):\Omega\to\mathbb{R}$ be $\mathcal{F}$-Borel-measurable with $E_{P}[|T|]<\infty$ or $T\geq 0$. Then Assumption 2.2 implies the following statements:
  • (i)
    $Q_{X}$ has a density $h$ with respect to $P_{X}$ which can be represented as
    $h(x)$ $\ =\ E_{P}[f\,|\,X=x],\quad P_{X}\text{-a.s.\ for}\ x\in\Omega_{X},$
    and it holds that
    $E_{Q}[T\,|\,X=x]$ $\ =\ \frac{E_{P}[T\,f\,|\,X=x]}{h(x)},$
    for $x\in\Omega_{X}$ with $h(x)>0$, i.e. with probability $1$ under $Q_{X}$.
  • (ii)
    $Q_{Y}$ has a density $g$ with respect to $P_{Y}$ which can be represented as
    $g(y)$ $\ =\ E_{P}[f\,|\,Y=y],\quad P_{Y}\text{-a.s.\ for}\ y\in\Omega_{Y},$
    and it holds that
    $E_{Q}[T\,|\,Y=y]$ $\ =\ \frac{E_{P}[T\,f\,|\,Y=y]}{g(y)},$
    for $y\in\Omega_{Y}$ with $g(y)>0$, i.e. with probability $1$ under $Q_{Y}$.
Proof of Proposition 2.3.
The statements about the densities follow from the observation that $\frac{dQ|\sigma(X)}{dP|\sigma(X)}=E_{P}[f\,|\,\sigma(X)]$ and $\frac{dQ|\sigma(Y)}{dP|\sigma(Y)}=E_{P}[f\,|\,\sigma(Y)]$ as well as (2.3).
The statements about the conditional expected values follow from the generalized Bayes formula, as stated in Lemma A1 of Tasche (2022), the formulae for the densities and (2.3). ∎
Since application of Fubini’s theorem (2.4a) cannot be justified in the absence of any assumptions on the properties of the source distribution $P$ in the setting as specified in Section 2.1, it is not possible to draw further conclusions about the composition of the density $f$ or the relationship between the posterior distributions of $X$ given $Y$ under the target distribution $Q$ and the source distribution $P$ respectively. This is in contrast to the situation studied by Tasche (2022). In Tasche (2022), thanks to the assumption that the label variable is categorical, more refined formulae for $f$ (Theorem 1) and the posterior probabilities (Corollary 2, a more general version of Eq. (2.4) of Saerens et al., 2002, but of similar shape) could be derived.
Section 2.4 specifies additional assumptions on the properties of the source distribution $P$. Based on those assumptions, we present follow-ups of the statements in Proposition 2.3 which facilitate numerical evaluations of the statistics related to the target distribution $Q$ among other things.
2.4 Assumptions on the source distribution and conditional densities
With additional assumptions on the properties of the source distribution $P$, the density $f$ according to Assumption 2.2 can be decomposed into ‘building blocks’ which allow for interpretations in terms of causality or chronological order and facilitate the demarcation of specific types of distribution shifts.
Assumption 2.4 (Regular conditional distributions).
The following two assumptions relating to the source distribution $P$ are to be understood in the setting of Section 2.1:
  • (i)
    There exists a conditional distribution $P_{Y|X}$ of $Y$ with respect to $X$ in the sense of Definition 2.1.
  • (ii)
    There exists a conditional distribution $P_{X|Y}$ of $X$ with respect to $Y$ in the sense of Definition 2.1.
Thanks to (2.4b) Assumption 2.4 immediately implies the following version of Proposition 2.3.
Corollary 2.5.
In the setting of Section 2.1, let $T=T(X,Y):\Omega\to\mathbb{R}$ be $\mathcal{F}$-Borel-measurable with $E_{P}[|T|]<\infty$ or $T\geq 0$.
  • (i)
    Under Assumptions 2.2 and 2.4 (i), the density $h$ of $Q_{X}$ with respect to $P_{X}$ can be represented as
    $h(x)$ $\ =\ \int f(x,y)\,P_{Y|X=x}(dy),\quad P_{X}\text{-a.s.\ for}\ x\in\Omega_{X},$
    and it holds that
    $E_{Q}[T\,|\,X=x]$ $\ =\ \frac{\int T(x,y)\,f(x,y)\,P_{Y|X=x}(dy)}{h(x)},\quad P_{Y}\text{-a.s.\ for}\ y\in\Omega_{Y},$
    for $x\in\Omega_{X}$ with $h(x)>0$, i.e. with probability $1$ under $Q_{X}$.
  • (ii)
    Under Assumptions 2.2 and 2.4 (ii), the density $g$ of $Q_{Y}$ with respect to $P_{Y}$ can be represented as
    $g(y)$ $\ =\ \int f(x,y)\,P_{X|Y=y}(dx),\quad P_{Y}\text{-a.s.\ for}\ y\in\Omega_{Y},$
    and it holds that
    $E_{Q}[T\,|\,Y=y]$ $\ =\ \frac{\int T(x,y)\,f(x,y)\,P_{X|Y=y}(dx)}{g(y)},$
    for $y\in\Omega_{Y}$ with $g(y)>0$, i.e. with probability $1$ under $Q_{Y}$.
Assumption 2.4 is also useful because of the following result that clarifies the concept of conditional density as considered in the distribution shift context of this paper.
Theorem 2.6.
  • (i)
    Under Assumptions 2.2 and 2.4 (i), let $h$ be a density of $Q_{X}$ with respect to $P_{X}$ and define $q_{Y|X}:\Omega_{X}\times\Omega_{Y}\to[0,\infty)$ by $q_{Y|X=x}(y)=\begin{cases}\frac{f(x,y)}{h(x)},&h(x)>0\\ 0,&h(x)=0\end{cases}$ for $x\in\Omega_{X}$ and $y\in\Omega_{Y}$. Then for $x\in\Omega$ with $h(x)>0$ (and therefore $Q_{X}$-a.s.) $q_{Y|X=x}$ is a density of $Q_{Y|X=x}$ with respect to $P_{Y|X=x}$.
  • (ii)
    Under Assumptions 2.2 and 2.4 (ii), let $g$ be a density of $Q_{Y}$ with respect to $P_{Y}$ and define $q_{X|Y}:\Omega_{X}\times\Omega_{Y}\to[0,\infty)$ by $q_{X|Y=y}(x)=\begin{cases}\frac{f(x,y)}{g(y)},&g(y)>0\\ 0,&g(y)=0\end{cases}$ for $x\in\Omega_{X}$ and $y\in\Omega_{Y}$. Then for $y\in\Omega$ with $g(y)>0$ (and therefore $Q_{Y}$-a.s.) $q_{X|Y=y}$ is a density of $Q_{X|Y=y}$ with respect to $P_{X|Y=y}$.
Proof of Theorem 2.6.
Observe that $(x,y)\mapsto q_{Y|X=x}(y)$ is $\mathcal{H}\otimes\mathcal{G}$-Borel measurable. Define $K:\Omega_{X}\times\mathcal{G}\to[0,\infty)$ by
[…]
with a fixed arbitrary $y_{0}\in\Omega_{Y}$, for $x\in\Omega_{X}$ and $G\in\mathcal{G}$. Then $K$ satisfies (i) of Definition 2.1 and $K(x,\cdot)$ is a measure on $(\Omega_{Y},\mathcal{G})$. By Fubini’s theorem (2.4a), it follows for $H\in\mathcal{H}$ and $G\in\mathcal{G}$ that
$Q[X\in H,Y\in G]$ $=\int_{H\cap\{h>0\}}\left(\int_{G}q_{Y|X=x}(y)\,P_{Y|X=x}(dy)\right)h(x)\,P_{X}(dx)$
$=\int_{H\cap\{h>0\}}K(x,G)\,Q_{X}(dx)$
$=\int_{H}K(x,G)\,Q_{X}(dx).$
By (2.2), this observation implies (iii) of Definition 2.1 for every $G\in\mathcal{G}$. That $K(x,\cdot)$ is a probability measure follows from (iii) of Definition 2.1 for $G=\Omega_{Y}$. Hence $Q_{Y|X}=K$ is a regular conditional distribution of $Y$ with respect to $X$ under $Q$. The density property of $q_{Y|X=x}$ for $x$ with $h(x)>0$ then follows from the definition of $K$. This implies (i) of Theorem 2.6. (ii) follows from (i) since the roles of $X$ and $Y$ may be swapped thanks to the symmetry of the setting in Section 2.1. ∎
In the following, the densities $q_{X|Y}$ and $q_{Y|X}$ as defined in Theorem 2.6 both are called conditional density. Based on Thorem 2.6, versions for general label spaces can be formulated of the ‘normal form’ of the joint density $f$ (Theorem 1, Tasche, 2022), the feature density $h$ (Corollary 1, Tasche, 2022), and of the ‘posterior correction formula’ [^3] Called ‘Bayes update rule’ in Donyavi et al. (2025). (Tasche, 2022, Corollary 2).
Corollary 2.7.
  • (i)
    Under Assumption 2.2 and Assumption 2.4 (ii), the density $f$ of the joint target distribution $Q$ of $X$ and $Y$ with respect to the joint source distribution $P$ can be represented in terms of the conditional density $q_{X|Y}$ as defined in Thorem 2.6 and the label density $g$ as
    $f(x,y)\ =\ q_{X|Y=y}(x)\,g(y),\quad\text{for}\ (x,y)\in\Omega.$ (2.6a)
  • (ii)
    Under Assumption 2.2 and Assumptions 2.4 (i) and (ii), the feature density $h$ of $Q_{X}$ with respect to $P_{X}$ can be written in terms of the conditional density $q_{X|Y}$, the label density $g$ and the source conditional distribution $P_{Y|X}$ as
[…]
  • (iii)
    Under Assumption 2.2 and Assumptions 2.4 (i) and (ii), the posterior probabilities of $Y$ given $X$ under $Q$ can be described in terms of the conditional density $q_{X|Y}$, the label density $g$ and the source conditional distribution $P_{Y|X}$ for all $x\in\Omega_{X}$ with $h(x)>0$ (i.e. $Q_{x}$-a.s.) by
    $Q_{Y|X=x}[G]\ =\ \frac{\int_{G}q_{X|Y=y}(x)\,g(y)\,P_{Y|X=x}(dy)}{h(x)},\quad\text{for}\ G\in\mathcal{G}.$ (2.6c)
Proof of Corollary 2.7..
(i) is basically the definition of $q_{X|Y=y}(x)$ in Thorem 2.6 (ii). (ii) is the combination of (i) and Corollary 2.5 (i). (iii) is implied also by (i) and Corollary 2.5 (i). ∎
2.5 Facilitating calculations by further assumptions
The setting of Tasche (2022) differs from the setting in Section 2.1 through the additional assumption of the label variable $Y$ being categorical. This is stated formally in the following assumption where without loss of generality the categories are supposed to be natural numbers.
Assumption 2.8 (Categorical label space).
In the setting of Section 2.1, the label space $\Omega_{Y}$ is finite with $\Omega_{Y}=\{1,\ldots,d\}$, $d\geq 2$. The $\sigma$-algebra $\mathcal{G}$ is the power set of $\Omega_{Y}$, i.e. $\mathcal{G}=\mathfrak{P}(\Omega_{Y})$.
Under Assumption 2.8, suppose in addition $0<P[Y=i]$, $i=1,\ldots,d$. Then thanks to the fact that $\Omega_{Y}$ is finite, Assumptions 2.4 (i) and (ii) both hold true.
Under distribution shift as expressed through Assumption 2.2, to reconcile Corollary 2.7 with the corresponding statements of Tasche (2022) under Assumption 2.8, observe that $q_{X|Y=i}=h_{i}$ for the conditional densities defined in Lemma 1 of Tasche (2022) and $g(i)=\frac{Q[Y=i]}{P[Y=i]}$ for $i=1,\ldots,d$. Then it follows from Corollary 2.7 (iii) with $G=\{i\}$ that
[…]
(2.7) restates Corollary 2 of Tasche (2022) in the notation of this paper. Similarly, Theorem 1 and Corollary 1 of Tasche (2022) are implied by Corollary 2.7 (i) and (ii) respectively.
In the case where the label variable $Y$ is not categorical, even if Assumption 2.4 (i) or Assumption 2.4 (ii) can be shown to be satisfied thanks to certain properties of the feature or label spaces (like in the case where $Y$ is real-valued, cf. the comments after Definition 2.1), it is not yet clear how $P_{Y|X}$ and $P_{X|Y}$ can be determined. For this reason the subsequent Assumption 2.9 is useful in so far as it implies both Assumption 2.4 (i) and Assumption 2.4 (ii) as well as convenient representations of $P_{Y|X}$ and $P_{X|Y}$ in Propostion 2.10 below.
Define in the setting of Section 2.1 the probability measure $P_{X}\otimes P_{Y}$ on $(\Omega,\mathcal{F})$ as the product measure of $P_{X}$ and $P_{Y}$. Then $P_{X}\otimes P_{Y}$ is uniquely determined as the distribution on $(\Omega,\mathcal{F})$ such that $X$ and $Y$ are independent, i.e. by the property
$P_{X}\otimes P_{Y}[H\times G]=P_{X}\otimes P_{Y}[X\in H,Y\in G]=P_{X}[H]\,P_{Y}[G],\quad H\in\mathcal{H},G\in\mathcal{G}.$
In general, it holds that $P_{X}\otimes P_{Y}\neq P$ but the marginal distributions of $X$ and $Y$ under $P_{X}\otimes P_{Y}$ and $P$ are equal.
Assumption 2.9 (Absolute continuity of source distribution).
[…]
$\frac{dP}{dP_{X}\otimes P_{Y}}\ =\ \varphi.$ (2.8)
Observe that Assumption 2.8 implies Assumption 2.9 if $0<P[Y=i]$, $i=1,\ldots,d$, with
$\varphi(x,i)=\frac{P[Y=i\,|\,X=x]}{P[Y=i]},\quad x\in\Omega_{X},i\in\Omega_{Y}=\{1,\ldots,d\}.$ (2.9)
The choice of $P_{X}\otimes P_{Y}$ as reference measure in Assumption 2.9 is convenient because no additional assumptions on the structural properties of the instance space $(\Omega,\mathcal{F})$ are required, in contrast e.g. to the case where the Lebesgue measure serves as reference measure. Nonetheless, having a density of $P$ with respect to $P_{X}\otimes P_{Y}$ does not exclude representations of $P$ with other densities for computational purposes. For instance, if $P_{X}$ and $P_{Y}$ have both Lebesgue densities, then by the density chain rule $P$ also has a Lebesgue density.
Intuitively, Assumption 2.9 is true if the supports of $P$ and $P_{X}\otimes P_{Y}$ have the same dimension. As Filipović and Schneider (2025, Section 5.2.2) point out, Assumption 2.9 is violated in particular if $X$ is a function of $Y$ or vice versa. Filipović and Schneider (2025, Lemma 5.3) provide a sufficient criterion for Assumption 2.9 to be true. Moreover, there are feasible methods for finding the density of $P$ with respect to $P_{X}\otimes P_{Y}$. Filipović and Schneider (2025, Section 5.2) propose to use a reproducing kernel Hilbert space (RKHS) method for estimating $\varphi$. They also comment on other methods for doing so from the previous literature on the subject. In addition, $\varphi$ can be determined by means of a binary probabilistic classifier, see Appendix A below. An advantage of the classification approach is that in principle the spaces $(\Omega_{X},\mathcal{H})$ and $(\Omega_{Y},\mathcal{G})$ need not satisfy any specific conditions for the approach to work.
Proposition 2.10.
Assumption 2.9 implies both Assumption 2.4 (i) and Assumption 2.4 (ii), with
$\begin{split}P_{Y|X=x}[G]&=\int_{G}\varphi(x,y)\,P_{Y}(dy),\quad\text{for}\ x\in\Omega_{X},G\in\mathcal{G},\quad\text{and}\\ P_{X|Y=y}[H]&=\int_{H}\varphi(x,y)\,P_{X}(dx),\quad\text{for}\ y\in\Omega_{Y},H\in\mathcal{H}.\end{split}$
Proof of Proposition 2.10.
Actually, Proposition 2.10 is a special case of Theorem 2.6. For instance, to prove the statement regarding Assumption 2.4 (i), substitute $P$ for $Q$, $P_{X}\otimes P_{Y}$ for $P$ and $\varphi$ for $f$ in Theorem 2.6. The theorem is applicable because under $P_{X}\otimes P_{Y}$ the marginal distributions of $X$ and $Y$ are regular conditional distributions of $X$ given $Y$ and $Y$ given $X$ respectively. Moreover, since the marginal distributions of $X$ and $Y$ are the same under $P$ and $P_{X}\otimes P_{Y}$, it follows that
$\begin{split}\frac{dP_{X}}{d(P_{X}\otimes P_{Y})_{X}}&=\frac{dP_{X}}{dP_{X}}=1\quad\text{and}\\ \frac{dP_{Y}}{d(P_{X}\otimes P_{Y})_{Y}}&=\frac{dP_{Y}}{dP_{Y}}=1,\end{split}$ (2.10)
and hence in the notation of Theorem 2.6 $g(y)=\int\varphi(x,y)\,P_{X}(dx)=1=\int\varphi(x,y)\,P_{Y}(dy)=h(x)$ $P_{Y}$-a.s. for $y\in\Omega_{Y}$ and $P_{X}$-a.s. for $x\in\Omega_{X}$. ∎
Thanks to Proposition 2.10, under Assumption 2.9 Fubini’s theorem for conditional distributions (2.4a) and the formulae (2.4b) for conditional expected values under $P$ obtain simpler shapes. For instance, the expected value of a $\mathcal{F}$-Borel measurable random variable $T:\Omega\to\mathbb{R},(x,y)\mapsto T(x,y)$ with $E_{P}[\lvert T(X,Y)\rvert$$\infty$ or $T\geq 0$ conditional on $X=x$ can be represented as
$E_{P}[T(X,Y)\,|\,X=x]=\int T(x,y)\,\varphi(x,y)P_{Y}(dy),\quad P_{X}\text{-a.s.~for}\ x\in\Omega_{X}.$ (2.11)
The same comment applies to the formulae from Corollary 2.5. The definitions of the conditional densities which are introduced in Theorem 2.6 remain unchanged under Assumption 2.9 such that for example there is no change in (2.6a) either. However, under Assumption 2.9 the density $h$ of $Q_{X}$ with respect to $P_{X}$ as represented in (2.6b) may be calculated via
[…]
Under Assumption 2.9, moreover the posterior correction formula (2.6c) can be rewritten (for $x\in\Omega_{X}$ with $h(x)>0$) as
$Q[Y\in G\,|\,X=x]\ =\ \frac{\int_{G}q_{Y|X=x}(y)\,g(y)\,\varphi(x,y)\,P_{Y}(dy)}{h(x)},\quad\ G\in\mathcal{G},$ (2.12b)
with $h(x)$ as in (2.12a).
3 Factorizable joint shift (FJS) for general label spaces
He et al. (2021) proposed FJS “to handle the co-existence of sampling bias in covariates and labels”, and consequently as a notion generalising both covariate and label shift. Tasche (2022, Remark 1) characterised FJS in the case of categorical label variables as invariance between the source and the target distribution up to constant factors of ratios of the feature densities conditional on different labels. In addition, he noted that in domain adaptation when only the feature marginal distribution of the target distribution is known (i.e. Problem (B) of Section 2), assuming FJS without constraints between the source and the target distributions does not uniquely determine the distribution shift. More recently, Dong et al. (2025) observed that FJS may be described as combination of a label and a covariate shift. This section generalises these observations for the case of FJS in the face of a general label space. In addition, solutions to Problems (A) and (B) of Section 2 under the assumption of FJS are discussed.
3.1 Characterising FJS
Definition 3.1.
In the setting of Section 2.1, $P$ and $Q$ are related through factorizable joint shift (FJS) if there are an $\mathcal{H}$-measurable function $\overline{h}:\Omega_{X}\to[0,\infty)$ and a $\mathcal{G}$-measurable function $\overline{g}:\Omega_{Y}\to[0,\infty)$ such that $\overline{h}(X)\,\overline{g}(Y)$ is a density of $Q$ with respect to $P$.
If $P$ and $Q$ are related through FJS then in particular Asssumption 2.2 is satisfied. Note that $\overline{h}$ and $\overline{g}$ in Definition 3.1 are unique only up to a constant factor $c>0$, in the sense that
$\frac{dQ}{dP}\quad=\quad\overline{h}_{c}(X)\,\overline{g}_{c}(Y),$ (3.1)
for $\overline{h}_{c}=c\,\overline{h}$ and $\overline{g}_{c}=\overline{g}/c$. As mentioned above, FJS encompasses both covariate shift and label shift as special cases.
Definition 3.2.
In the setting of Section 2.1, the notions of covariate and label shift respectively are defined as follows:
  • (i)
    $Q$ and $P$ are related through covariate shift if for each $G\in\mathcal{G}$ there is an $\mathcal{H}$-Borel measurable function $\eta_{G}:\Omega_{X}\to[0,1]$ such that $Q[Y\in G\,|\,X=x]=\eta_{G}(x)$ $Q_{X}$-a.s. and $P[Y\in G\,|\,X=x]=\eta_{G}(x)$ $P_{X}$-a.s. hold.
  • (ii)
    $Q$ and $P$ are related through label shift if for each $H\in\mathcal{H}$ there is a $\mathcal{G}$-Borel measurable function $\gamma_{H}:\Omega_{y}\to[0,1]$ such that $Q[X\in H\,|\,Y=y]=\gamma_{H}(y)$ $Q_{Y}$-a.s. and $P[X\in H\,|\,Y=y]=\gamma_{H}(y)$ $P_{Y}$-a.s. hold.
Proposition 3.3.
In the setting of Section 2.1 and under Assumption 2.2, both covariate shift and label shift in the sense of Definition 3.2 imply FJS in the sense of Definition 3.1 between the source distribution $P$ and the target distribution $Q$.
Conversely, if $P$ and $Q$ are related through FJS with $\overline{g}=1$ then $P$ and $Q$ are related also through covariate shift. If $P$ and $Q$ are related through FJS with $\overline{h}=1$ then $P$ and $Q$ are related also through label shift.
Proof of Proposition 3.3..
Denote by $f$ the density of $Q$ with respect to $P$ that exists by Assumption 2.2. Define the $\mathcal{H}$-measurable function $h\geq 0$ by $h(x)=E_{P}[f\,|\,X=x]$ for $x\in\Omega_{X}$. Then for $H\in\mathcal{H}$ and $G\in\mathcal{G}$, under covariate shift it follows that
$Q[X\in H,Y\in G]$ $=E_{Q}[\mathbf{1}_{\{X\in H\}}\,\eta_{G}(X)]$
$=E_{P}[f\,\mathbf{1}_{\{X\in H\}}\,\eta_{G}(X)]$
$=E_{P}\bigl[E_{P}[f\,|\,\sigma(X)]\,\mathbf{1}_{\{X\in H\}}\,P[Y\in G\,|\,\sigma(X)\bigr]$
$=E_{P}[h(X)\,\mathbf{1}_{\{X\in H,Y\in G\}}].$
By (2.1), $\bigl\{\{X\in H,Y\in G\}:H\in\mathcal{H},G\in\mathcal{G}\bigr\}$ is a $\cap$-stable generator of $\mathcal{F}$, hence by the measure uniqueness theorem (Bauer, 1981, Theorem 1.5.5) $h(X)$ is a density of $Q$ with respect to $P$ on $(\Omega,\mathcal{F})$. Thus, $P$ and $Q$ are related through FJS with $\overline{h}=h$ and $\overline{g}=1$. Thanks to the symmetry of the setting with respect to $X$ and $Y$, we also have shown that label shift implies FJS with $\overline{h}=1$ and $\overline{g}=E_{P}[f\,|\,Y=\cdot]$.
The converse claims follow from Lemma 4.2 of Tasche (2024). ∎
By the density chain rule, a sequence of FJS results in another FJS. In particular, by Proposition 3.3 any sequence of covariate and label shifts gives rise to an FJS. Conversely, any FJS can be interpreted as consecutive label and covariate shifts or consecutive covariate and label shifts. The following proposition provides a streamlined version of Theorem 1 of Dong et al. (2025) in the more general setting of Section 2.1.
Proposition 3.4.
Let $P$ and $Q$ are related through FJS in the sense of Definition 3.1. Define probability measures $Q_{L}$ and $Q_{C}$ on $(\Omega,\mathcal{F})$ by
$\frac{dQ_{L}}{dP}$ $=\frac{\overline{g}(Y)}{E_{P}\bigl[\overline{g}(Y)\bigr]},$ (3.2a)
$\frac{dQ_{C}}{dQ_{L}}$ $=\frac{\overline{h}(X)}{E_{Q_{L}}\bigl[\overline{h}(X)\bigr]}.$ (3.2b)
Then the following statements hold true:
  • (i)
    $Q_{L}$ and $Q_{C}$ are well-defined since it holds that $E_{P}\bigl[\overline{g}(Y)\bigr]>0$ and $E_{Q_{L}}\bigl[\overline{h}(X)\bigr]>0$.
  • (ii)
    $Q=Q_{C}$.
  • (iii)
    For all $F\in\mathcal{F}$ it holds that $Q_{L}[F\,|\,\sigma(Y)]=P[F\,|\,\sigma(Y)]$ and $Q_{C}[F\,|\,\sigma(X)]\ =\ Q[F\,|\,\sigma(X)]$. In other words, $P$ and $Q_{L}$ are related through label shift while $Q_{C}$ and $Q$ are related through covariate shift.
Proof of Proposition 3.4.
  • (i)
    $E_{P}\bigl[\overline{g}(Y)\bigr]=0$ would imply $\overline{g}(Y)=0$ $P$-a.s. and therefore also $E_{P}\bigl[\overline{h}(X)\,\overline{g}(Y)\bigr]=0$. This would contradict $E_{P}\bigl[\overline{h}(X)\,\overline{g}(Y)\bigr]=1$ which is true by definition of FJS. Thus $E_{P}\bigl[\overline{g}(Y)\bigr]>0$ is true. From the definition of $Q_{L}$ it follows that
    $E_{Q_{L}}\bigl[\overline{h}(X)\bigr]\ =\ \frac{E_{P}\bigl[\overline{h}(X)\,\overline{g}(Y)\bigr]}{E_{P}\bigl[\overline{g}(Y)\bigr]}\ =\ \frac{1}{E_{P}\bigl[\overline{g}(Y)\bigr]}\ >\ 0.$
    Hence also $Q_{C}$ is well-defined.
  • (ii)
    For all $F\in\mathcal{F}$ it follows from the definitions of $Q_{C}$ and $Q_{L}$ that
    $Q_{C}[F]\ =\ \frac{E_{Q_{L}}\bigl[\overline{h}(X)\,\mathbf{1}_{F}\bigr]}{E_{Q_{L}}\bigl[\overline{h}(X)\bigr]}\ =\ \,E_{P}\bigl[\overline{h}(X)\,\overline{g}(Y)\,\mathbf{1}_{F}\bigr]\frac{E_{P}\bigl[\overline{g}(Y)\bigr]}{E_{P}\bigl[\overline{g}(Y)\bigr]}\ =\ Q[F].$
  • (iii)
    Both equations in (iii) follow from Proposition 3.3. ∎
Against the background of the description of FJS as a sequence of covariate and label shifts, the question arises how much FJS is constrained when compared to non-specific joint distribution shift between $Q$ and $P$. The following result helps to answer this question by presenting the relations between the functions $\overline{h}$ and $\overline{g}$ which characterise FJS on the one hand and the feature density $h$ and the label density $g$ on the other hand.
Theorem 3.5.
Assume the setting of Section 2.1.
  1. (i)
    Let $P$ and $Q$ be related through FJS in the sense of Definition 3.1 such that there are an $\mathcal{H}$-measurable function $\overline{h}:\Omega_{X}\to[0,\infty)$ and a $\mathcal{G}$-measurable function $\overline{g}:\Omega_{Y}\to[0,\infty)$ with the property that $\overline{h}(X)\,\overline{g}(Y)$ is a density of $Q$ with respect to $P$. Denote by $h$ a density of $Q_{X}$ with respect to $P_{X}$ on $\mathcal{H}$ and by $g$ a density of $Q_{Y}$ with respect to $P_{Y}$ on $\mathcal{G}$. Then it follows that $P$-a.s.
    $h(X)=\overline{h}(X)\,E_{P}[\overline{g}(Y)\,|\,\sigma(X)]\qquad\text{and}\qquad g(Y)=\overline{g}(Y)\,E_{P}[\overline{h}(X)\,|\,\sigma(Y)].$ (3.3)
  2. (ii)
    Suppose that $h$ is a density of some probability measure $Q_{X}^{\ast}$ on $(\Omega_{X},\mathcal{H})$ with respect to $P_{X}$ and that $g$ is a density of some probability measure $Q_{Y}^{\ast}$ on $(\Omega_{Y},\mathcal{G})$ with respect to $P_{Y}$. In addition, let $\overline{h}:\Omega_{X}\to[0,\infty)$ be an $\mathcal{H}$-measurable function and $\overline{g}:\Omega_{Y}\to[0,\infty)$ be a $\mathcal{G}$-measurable function such that (3.3) holds true. Then $\overline{h}(X)\,\overline{g}(Y)$ is a probability density on $(\Omega,\mathcal{F})$, and for $Q$ defined by $\frac{dQ}{dP}=\overline{h}(X)\,\overline{g}(Y)$ it holds that $Q_{X}=Q_{X}^{\ast}$ and $Q_{Y}=Q_{Y}^{\ast}$.
Proof of Theorem 3.5.
  • (i)
    Assume that $\overline{h}(X)\,\overline{g}(Y)$ is a density of $Q$ with respect to $P$. Since $h(X)$ is a density of $Q|\sigma(X)$ with respect to $P|\sigma(X)$ it follows that
    $h(X)\ =\ E_{P}\bigl[\overline{h}(X)\,\overline{g}(Y)\,|\,\sigma(X)\bigr]\ =\ \overline{h}(X)\,E_{P}\bigl[\overline{g}(Y)\,|\,\sigma(X)\bigr].$
    The second equation in (3.3) follows similarly because $g(Y)$ is a density of $Q|\sigma(Y)$ with respect to $P|\sigma(Y)$.
  • (ii)
    Suppose that (3.3) is true for densities $h$ and $g$ and functions $\overline{h}$ and $\overline{g}$ as described in Theorem 3.5 (ii). Then it holds that
    $E_{P}[\overline{h}(X)\,\overline{g}(Y)]=E_{P}\bigl[\overline{h}(X)\,E_{P}[\overline{g}(Y)\,|\,\sigma(X)]\bigr]=E_{P}[h(X)]=1.$
    Hence $\frac{dQ}{dP}=\overline{h}(X)\,\overline{g}(Y)$ defines a probability measure $Q$ on $(\Omega,\mathcal{F})$. (3.3) implies that $h(X)$ is the marginal density of $Q$ with respect to $P$ on $\sigma(X)$ and $g(Y)$ is the marginal density of $Q$ with respect to $P$ on $\sigma(Y)$. From this observation it follows that $Q_{X}=Q^{\ast}_{X}$ on $(\Omega_{X},\mathcal{H})$ and $Q_{Y}=Q^{\ast}_{Y}$ on $(\Omega_{Y},\mathcal{G})$. ∎
As shown in Remark 3.8 below, Theorem 2 of Tasche (2022) is a special case of Theorem 3.5, for categorical label variables.
Remark 3.6.
Suppose that $Q$ and $P$ are related through FJS and that Assumption 2.4(i) is satisfied. Observe that then by (3.3) it holds that $P_{X}$-a.s.
$\{h>0\}\ =\ \{\overline{h}>0\}\cap\bigl\{E_{P}[\overline{g}(Y)\,|\,X=\cdot]>0\bigr\}.$
Therefore, it follows for $x\in\Omega_{X}$ with $h(x)>0$ that $\frac{\overline{h}(x)}{h(x)}=\frac{1}{E_{P}[\overline{g}(Y)\,|\,X=x]}$. Accordingly, under FJS the conditional density $q_{Y|X}$ as defined in Theorem 2.6 (i) can be represented for all $x$ with $h(x)>0$ and hence $Q_{X}$-a.s. as
$q_{Y|X=x}(y)\ =\ \frac{\overline{g}(y)}{\int\overline{g}(z)\,P_{Y|X=x}(dz)},\quad y\in\Omega_{Y}.$ (3.4a)
In the case where the label space $\Omega_{Y}$ is a subset of $\mathbb{R}$, (3.4a) implies the following formula for the expected value of $Y$ conditional on $X$ under the target distribution $Q$:
$E_{Q}[Y\,|\,X=x]\ =\ \frac{\int y\,\overline{g}(y)\,P_{Y|X=x}(dy)}{\int\overline{g}(z)\,P_{Y|X=x}(dz)},\quad Q_{X}\text{-a.s.\ for}\ x\in\Omega_{X}.$ (3.4b)
Moreover, (3.4a) suggests the following interpretation of the factors $\overline{h}$ and $\overline{g}$ in the definition of FJS:
  • In general, $\overline{g}$ does not represent a marginal density of the label variable $Y$ as in the case of label shift (i.e. when $\overline{h}=1$) but rather is characterised by being a conditional density of $Q_{Y|X}$ with respect to $P_{Y|X}$, after normalisation. Due to the symmetry of the setting in $X$ and $Y$, $\overline{h}$ is not a marginal density of the feature variable $X$ as in the case of covariate shift but rather is characterised by being a conditional density of $Q_{X|Y}$ with respect to $P_{X|Y}$, again after normalisation.
  • FJS is more homogeneous than non-specific distribution shift in so far as the target conditional densities of the labels given the features are all equal (up to normalisation). ∎
3.2 Solving Problems (A) and (B) by assuming FJS
How to solve (3.3) for $\overline{h}$ and $\overline{g}$ if marginal densities $h$ and $g$ are given, i.e. if Problem (A) of Section 2 ought to be solved? Corollary 3.7 disentangles (3.3) to arrive at an integral equation for $\overline{g}$ and a representation of $\overline{h}$ in terms of $h$ and $\overline{g}$.
Corollary 3.7.
  1. (i)
    Under the assumptions and with the notation of Theorem 3.5 (ii), suppose that the density $g$ is $P_{Y}$-a.s. positive. Then $P\bigl[E_{P}[\overline{g}(Y)\,|\,\sigma(X)]>0\bigr]=1$ follows and (3.3) is equivalent to
    $\overline{h}(X)=\frac{h(X)}{E_{P}[\overline{g}(Y)\,|\,\sigma(X)]}\qquad\text{and}\qquad g(Y)=\overline{g}(Y)\,E_{P}\left[\frac{h(X)}{E_{P}[\overline{g}(Y)\,|\,\sigma(X)]}\,\Big|\,\sigma(Y)\right],\quad P\text{-a.s.}$ (3.5a)
  2. (ii)
    In addition to the assumptions of (i), suppose that the source distribution $P$ on $(\Omega,\mathcal{F})$ satisfies Assumptions 2.4 (i) and (ii) and define $\psi(y)=\frac{\overline{g}(y)}{g(y)}$ for $y\in\Omega_{Y}$. Then (3.5a) and (3.3) are equivalent to
    $\overline{h}(x)$ $=\frac{h(x)}{\int\psi(z)\,g(z)\,P_{Y|X=x}(dz)},\ P_{X}\text{-a.s.~for}\ x\in\Omega_{X},\ \text{and}$ (3.5b)
    $1$ $=\psi(y)\int\frac{h(x)}{\int\psi(z)\,g(z)\,P_{Y|X=x}(dz)}\,P_{X|Y=y}(dx),\ P_{Y}\text{-a.s.~for}\ y\in\Omega_{Y}.$ (3.5c)
Proof of Corollary 3.7.
  • (i)
    (3.3) implies $\{g(Y)>0\}=\{\overline{g}(Y)>0\}\cap\bigl\{E_{P}[\overline{h}(X)\,|\,\sigma(Y)]>0\bigr\}\subset\{\overline{g}(Y)>0\}$. Hence it follows that $P[\overline{g}(Y)>0]=1$. Moreover, it holds that
    $E_{P}\bigl[\overline{g}(Y)\,\mathbf{1}_{\{E_{P}[\overline{g}(Y)\,|\,\sigma(X)]=0\}}\bigr]\ =\ E_{P}\bigl[E_{P}[\overline{g}(Y)\,|\,\sigma(X)]\,\mathbf{1}_{\{E_{P}[\overline{g}(Y)\,|\,\sigma(X)]=0\}}\bigr]\ =\ 0.$
    This observation implies $P\bigl[\overline{g}(Y)>0,E_{P}[\overline{g}(Y)\,|\,\sigma(X)]=0\bigr]=0$. Taking into account $P[\overline{g}(Y)>0]=1$, it follows that $P\bigl[E_{P}[\overline{g}(Y)\,|\,\sigma(X)]=0\bigr]=0$. Therefore, division by $E_{P}[\overline{g}(Y)\,|\,\sigma(X)]$ is possible and the equivalence of (3.3) and (3.5a) follows.
  • (ii)
    Follows from (i) and (2.4b).∎
Remark 3.8.
Under Assumption 2.8 ‘Categorical label space’, Corollary 3.7 can be reconciled with Theorem 2 of Tasche (2022) as follows:
  • Transformed into the notation of Corollary 3.7, (9a) of Tasche (2022) reads $b=\overline{g}(Y)$, with $\frac{Q[Y=y]}{P[Y=y]}=g(y)$ and $\varrho(y)=\frac{\psi(y)}{\psi(d)}$ for $y=1,\ldots,d-1$.
  • In the notation of Corollary 3.7, when taking into account that $g$ in Tasche (2022) stands for $\overline{h}(X)$, $h$ for $h(X)$ and $\int\psi(z)\,g(z)\,P_{Y|X=x}(dz)=\sum_{i=1}^{d}\psi(i)\,g(i)\,P[Y=i\,|\,X=x]$, (9b) of Tasche (2022) turns out to be (3.5b).
  • (9c) of Tasche (2022) is mutatis mutandis nothing else but (3.5c).∎
Remark 3.9.
In the setting of Section 2.1, suppose that Assumption 2.4 (ii) is satisfied and that $Q$ and $P$ are related through FJS with $\frac{dQ}{dP}=\overline{h}(X)\,\overline{g}(Y)>0$ $P$-a.s.
Then it follows that $P$-a.s. $\overline{h}(X)>0$ and $\overline{g}(Y)>0$ and consequently also $E_{P}[\overline{h}(X)\,|\,\sigma(Y)]>0$ and $E_{P}[\overline{g}(Y)\,|\,\sigma(X)]>0$. As a further consequence, by Theorem 3.5 (i), also the feature density $h(X)$ and the label density $g(Y)$ are $P$-a.s. positive.
Theorem 2.6 (ii) in turn implies $P$-a.s. for $(x,y_{0}),(x,y_{1})\in\Omega$ that
$\frac{q_{X|Y=y_{1}}(x)}{q_{X|Y=y_{0}}(x)}\ =\ \frac{g(y_{0})}{g(y_{1})}\,\frac{\overline{g}(y_{1})}{\overline{g}(y_{0})}\ =\ \frac{\psi(y_{1})}{\psi(y_{0})},$ (3.6)
with $\psi=\frac{\overline{g}}{g}$ as in Corollary 3.7 (ii). By (3.6), the ratio of the target conditional feature densities evaluated at the same feature $x$ but for two different labels $y_{0}$ and $y_{1}$ does not depend on the value of the feature. In other words, $q_{X|Y=y_{1}}$ and $q_{X|Y=y_{0}}$ differ only by a calibration factor and are even identical in the case $\overline{g}=g$ (or $\psi=1$).
Based on a similar observation, Tasche (2022, Remark 1) suggested talking about “scaled density ratios” shift instead of FJS. ∎
Remark 3.10.
With the notation of Corollary 3.7 (ii), under Assumption 2.4 (i) (3.4a) of Remark 3.6 implies the posterior correction formula for general label spaces under FJS:
$Q_{Y|X=x}[G]\ =\ \frac{\int_{G}\psi(y)\,g(y)\,P_{Y|X=x}(dy)}{\int_{\Omega_{Y}}\psi(y)\,g(y)\,P_{Y|X=x}(dy)},\quad G\in\mathcal{G},\ Q_{X}\text{-a.s.\ for}\ x\in\Omega_{X}.$ (3.7)
(3.7) also follows from combining (2.6c) with (3.6). Eq. (12) of Tasche (2022) is the special case of (3.7) for categorical label spaces with $G=\{j\}$.
3.2.1 Solving Problem (A) by assuming FJS
Under the assumption of FJS, an ansatz for solving Problem (A) as phrased in Section 2 could be the following:
  • In Problem (A) the marginal distributions of $X$ and $Y$ under $Q$ are known (in addition to the source distribution $P$ including the conditional distributions $P_{X|Y}$ and $P_{Y|X}$). Hence solve the integral equation (3.5c) for the remaining unknown $\psi$. If the label space is categorical with $\Omega_{Y}=\{1,\ldots,d\}$ then (3.5c) becomes a (d1)(d-1)-dimensional non-linear equation system for $\varrho(y)=\frac{\psi(y)}{\psi(d)}$, $y=1,\ldots,d-1$ (Tasche, 2022).
  • Once $\psi$ has been determined, use $\overline{g}=\psi\,g$ and (3.5b) to infer the density $f(x,y)=\overline{h}(x)\,\overline{g}(y)$ of $Q$ with respect to $P$. Alternatively, (3.4a) can be used to infer the conditional density $q_{Y|X}$ such that predictions like (3.4b) can be made.
Under Assumption 2.9 ‘Absolute continuity of source distribution’, (3.5c) can be simplified thanks to Proposition 2.10:
$\begin{split}1&=\psi(y)\int\frac{h(x)\,\varphi(x,y)}{\int\psi(z)\,g(z)\,\varphi(x,z)\,P_{Y}(dz)}\,P_{X}(dx)\\ &=\psi(y)\int\frac{\varphi(x,y)}{\int\psi(z)\,\varphi(x,z)\,Q_{Y}(dz)}\,Q_{X}(dx).\end{split}$ (3.8)
On the one hand (3.8) is useful because the integrals with respect to conditional distributions have been replaced with integrals with respect to less complex marginal distributions. On the other hand, (3.8) (thanks to the second row) allows to make use of observations of the target feature variable and the target label variable if the integrals are approximated by sample means.
Moreover, for known marginal distributions $Q_{X}$ and $Q_{Y}$ as in Problem (A), (3.8) suggests the following interative approach to determining the function $\psi$:
  1. 1)
    Let $\psi_{0}(y)=1$ for all $y\in\Omega_{Y}$, or choose any other initialisation which does not hinder convergence of the iteration.
  2. 2)
    With known $\psi_{n}$ calculate $\psi_{n+1}$ as
    $\psi_{n+1}(y)\ =\ \left(\int\frac{\varphi(x,y)}{\int\psi_{n}(z)\,\varphi(x,z)\,Q_{Y}(dz)}\,Q_{X}(dx)\right)^{-1},\quad y\in\Omega_{Y}.$ (3.9)
  3. 3)
    Repeat 2) for $n=0,1,\ldots$ until an appropriate stop criterion is fulfilled.
Note that the existence of $\lim_{n\to\infty}\psi_{n}(y)=\psi^{\ast}(y)$ for all $y$ of the sequence defined by (3.9) does not necessarily always imply that $\psi^{\ast}$ fulfills (3.8) because without further conditions it cannot be excluded that
$\lim_{n\to\infty}\int\frac{\varphi(x,y)}{\int\psi_{n}(z)\,\varphi(x,z)\,Q_{Y}(dz)}\,Q_{X}(dx)\ \neq\ \int\frac{\varphi(x,y)}{\int\lim_{n\to\infty}\psi_{n}(z)\,\varphi(x,z)\,Q_{Y}(dz)}\,Q_{X}(dx).$ (3.10)
Clearly, (3.10) is not a problem in the case of a categorical label space. Tasche (2022, Proposition 2) proved existence and uniqueness of a solution to (3.8) for the case of binary labels.
3.2.2 Solving Problem (B) by assuming covariate shift
A straightforward way to solving Problem (B) as formulated in Section 2 is by assuming that the target distribution $Q$ and the source distribution $P$ are related through covariate shift in the sense of Definition 3.2. Covariate shift is implied by “missing at random” sample selection bias (Moreno-Torres et al., 2012, Section 6.1) such that there are applications where covariate shift is the only reasonable assumption.
Problem (B) means that a target feature distribution $Q^{\ast}_{X}$ is pre-specified by its density $h$ with respect to the source feature distribution $P_{X}$. According to Proposition 3.3, then letting $\frac{dQ}{dP}=h(X)$ defines a joint distribution $Q$ of features $X$ and labels $Y$ such that $Q$ and $P$ are related through covariate shift and it holds that $Q_{X}=Q^{\ast}_{X}$. Thus, Problem (B) is solved in principle.
However, Storkey (2009, Section 5.1) commented that $\ldots$ there is some benefit to be obtained by doing some-thing different in the case of covariate shift. The argument here is that these papers indicate a computational benefit rather than a fundamental modelling benefit.” The computational aspects of covariate shift have been extensively studied in the literature and are not discussed in this paper. See for example Kouw and Loog (2021, Section 3.1) for a review of the literature on domain adaptation under covariate shift.
3.2.3 Solving Problem (B) by assuming label shift or constrained FJS
Recall that in Problem (B) as phrased in Section 2 only the feature marginal distribution $Q_{X}$ (or its density $h$ with respect to $P_{X}$) is known of the source distribution $Q$. While most of the time the aim in Problem (B) is to come up with an estimate of the target posterior distribution $Q_{Y|X}$ of the labels conditional on the features, thanks to the law of total probability the marginal label distribution $Q_{Y}$ then is a byproduct. In the literature, the quest for $Q_{Y}$ in the case of categorical labels is called ‘quantification’ ‘counting’, ‘class probability re-estimation’, ‘class prior estimation’ or ‘class distribution estimation’ (Esuli et al., 2023, p. 2). In the case of general normed label spaces, Azizzadenesheli (2022) referred to the estimation of $Q_{Y}$ as ‘Importance Weight Estimation’.
Under the assumption of FJS, Corollary 3.7 and Theorem 3.5 (ii) together suggest the following ansatz for solving Problem (B).
  • To reduce the number of unknowns in (3.5c), choose $\psi$ constant with $\psi(y)=1$ for all $y\in\Omega_{Y}$. This choice is motivated by the observation that in the case of a categorical label variable with $\Omega_{Y}=\{1,\ldots,d\}$, (3.5c) is equivalent to the maximum likelihood equation system for the target prior probabilities $Q[Y=y]$, $y=1,\ldots,d-1$ (Tasche, 2022, Section 4.1). Another motivation could be a GLS assumption, see Corollary 4.2 and the related comments in Section 4 below.
  • Solve the integral equation (3.5c) with $\psi=1$ for the function $g$ under the constraints $g\geq 0$ and $\int g(y)\,P_{Y}(dy)=1$ to obtain the density $g$ of the label distribution $Q_{Y}$ with respect to $P_{Y}$. Again, it might be preferable to make use of the simplied equation (3.8) if possible.
  • Once $g$ has been determined, use $\overline{g}=g$ and (3.5b) to infer the density $f(x,y)=\overline{h}(x)\,\overline{g}(y)$ of $Q$ with respect to $P$. Alternatively, again (3.4a) can be used to infer the conditional density $q_{Y|X}$.
For the case of categorical labels, Tasche (2014, Theorem 1 (b)) provided a result on the uniqueness of the solution $g$ of (3.5c) with $\psi=1$. Peters and Coberly (1976) analysed the properties of a fixed point version of (3.5c) in the case of categorical labels and analysed an iterative solution approach.
(3.4a) together with Corollary 2.5 (ii) also is the basis for a ‘general label space’ variant of the expectation-maximisation (EM) algorithm for estimating target prior class probabilities (Saerens et al., 2002, Eq. (2.9)). This method for determining the target label density $g$ of $Q_{Y}$ with respect to $P_{Y}$ reads as follows:
  1. 1)
[…]
  2. 2)
    With $g_{n}$ known calculate
    $q_{Y|X=x}^{(n)}(y)\ =\ \frac{g_{n}(y)}{\int g_{n}(z)\,P_{Y|X=x}(dz)},\quad y\in\Omega_{Y},x\in\Omega_{X}.$ (3.11a)
  3. 3)
    With $q_{Y|X}^{(n)}$ known calculate
    $g_{n+1}(y)\ =\ \int q_{Y|X=x}^{(n)}(y)\,h(x)\,P_{X|Y=y}(dx),\quad y\in\Omega_{Y}.$ (3.11b)
  4. 4)
    Repeat 2) and 3) for $n=0,1,\ldots$ until an appropriate stop criterion is fulfilled.
Under Assumption 2.9 ‘Absolute continuity of source distribution’, (3.11a) and (3.11b) also can be simplified [^6] In order to exactly obtain (2.9) of Saerens et al. (2002) as a special case with categorical label space, define $q^{\ast}_{n}(x,y)=q_{Y|X=x}^{(n)}(y)\,\varphi(x,y)$ and rephrase (3.12) for $q^{\ast}_{n}(x,y)$ instead of $q_{Y|X=x}^{(n)}(y)$. thanks to Proposition 2.10:
$\begin{split}q_{Y|X=x}^{(n)}(y)&=\ \frac{g_{n}(y)}{\int g_{n}(z)\,\varphi(x,z)\,P_{Y}(dz)},\quad y\in\Omega_{Y},x\in\Omega_{X},\\ g_{n+1}(y)&=\ \int q_{Y|X=x}^{(n)}(y)\,\varphi(x,y)\,Q_{X}(dx),\quad y\in\Omega_{Y},\end{split}$ (3.12)
with $\varphi$ as defined in (2.8).
In the following analyses of the EM algorithm specified by (3.11a) and (3.11b), we make use of the standard notion of Kullback-Leibler divergence (cf. Section 1.6.1 of Bishop, 2006).
Definition 3.11.
Let $(\overline{\Omega},\mathcal{M},\mu)$ be a probability space. For $\mathcal{M}$-Borel measurable functions $\lambda_{0}\geq 0$ and $\lambda_{1}\geq 0$ with $\int\lambda_{i}\,d\mu=1$, $\mu[\lambda_{i}=0]=0$ for $i=0,1$ and $\int\lambda_{0}\left|\log\left(\frac{\lambda_{0}}{\lambda_{1}}\right)\right|d\mu<\infty$, the Kullback-Leibler divergence $\mathrm{KL}_{\mu}(\lambda_{0}\parallel\lambda_{1})$ of $\lambda_{1}$ with respect to $\lambda_{0}$ is defined as $\mathrm{KL}_{\mu}(\lambda_{0}\parallel\lambda_{1})=\int\lambda_{0}\,\log\left(\frac{\lambda_{0}}{\lambda_{1}}\right)d\mu$.
Jensen’s inequality implies $\mathrm{KL}_{\mu}(\lambda_{0}\parallel\lambda_{1})\geq 0$ and $\mathrm{KL}_{\mu}(\lambda_{0}\parallel\lambda_{1})=0$ if and only if $\lambda_{0}=\lambda_{1}$ $\mu$-a.s. Accordingly, $\lambda_{0}$ and $\lambda_{1}$ are the more similar, the closer to naught $\mathrm{KL}_{\mu}(\lambda_{0}\parallel\lambda_{1})$ is. Minimising $\mathrm{KL}_{\mu}(\lambda_{0}\parallel\lambda_{1})$ by changing $\lambda_{1}$ may be interpreted as a population version of maximum likelihood estimation.
Theorem 3.12 generalises the appendix of Saerens et al. (2002) “Derivation of the EM Algorithm” from categorical label spaces to general label spaces. Thus, for non-categorical label spaces, Theorem 3.12 establishes the EM algorithm as a potential alternative to the ‘inverse operator’ approach (comparable to ‘adjusted classify & count’ in the case of categorical labels) by Azizzadenesheli (2022). Proposition 3.14 generalises Theorem 1 on ‘exact fit’ of Tasche (2014). Proofs for Theorem 3.12 and Proposition 3.14 are presented below in Appendix B.
Theorem 3.12.
Under Assumptions 2.4 (i) and (ii), let $(\Omega,\mathcal{F},P)=(\Omega_{X}\times\Omega_{Y},\sigma(X,Y),P_{(X,Y)})$ be a source distribution of features $X$ and labels $Y$ as defined in Section 2.1. Assume that there is a some feature distribution $Q^{\ast}_{X}$ with a density $h>0$ with respect to $P_{X}$. In addition [^7] A further implicit assumption for Theorem 3.12 is that all integrals used in (3.13) and the proof of the theorem exist and are finite. In particular, the assumptions $E_{P}\bigl[h(X)\,|\log(h(X)|\bigr]<\infty$ and $E_{P}\bigl[h(X)\,|\log(h_{n}(X)|\bigr]<\infty$ for $n=0,1,\ldots$ ought to be made for Theorem 3.12 (ii)., suppose there is an initial label density $g_{0}>0$ with respect to $P_{Y}$. Then the following statements hold true:
  • (i)
    All functions $q^{(n)}_{Y|X=x}$ for $x\in\Omega_{X}$ and $n=0,1,\ldots$ and $g_{n}$ for $n=1,2,\ldots$ defined recursively by (3.11a) and (3.11b) are positive and densities with respect to $P_{Y|X=x}$ and $P_{Y}$ respectively.
  • (ii)
    Define a sequence $Q^{(n)}$, $n=0,1,\ldots$, of target distributions on $(\Omega,\mathcal{F})$ by letting $\frac{dQ^{(n)}}{dP}=g_{n}(Y)$ such that by Proposition 3.3 all $Q^{(n)}$ are related to $P$ through label shift. With $h_{n}=\frac{dQ^{(n)}_{X}}{dP_{X}}=E_{P}[g_{n}(Y)\,|\,X=\cdot]$, then it holds that
    $\mathrm{KL}_{P_{X}}(h\parallel h_{n+1})\ \leq\ \mathrm{KL}_{P_{X}}(h\parallel h_{n}),\quad\text{for all}\ n=0,1,\ldots.$ (3.13)
Remark 3.13 (Interpretation of Theorem 3.12).
Theorem 3.12 describes the EM algorithm as specified by (3.11a) and (3.11b) as the construction of a sequence of target joint distributions $Q^{(n)}$ of features $X$ and labels $Y$ such that $Q^{(n)}$ and the source joint distibution $P$ are related through label shift. Each $Q^{(n)}$ entails a target feature distribution $Q^{(n)}_{X}$ with a $P_{X}$-density $h_{n}$. As a result, there is an accompanying sequence of Kullback Leibler divergences $\mathrm{KL}_{P_{X}}(h\parallel h_{n})$ of $Q^{(n)}_{X}$ with respect to the pre-specified feature distribution $Q^{\ast}_{X}$ where both feature distributions are represented through their respective densities with respect to the source feature distribution $P_{X}$. By (3.13), this sequence decreases such that $Q^{(n+1)}_{X}$ resembles $Q^{\ast}_{X}$ more than $Q^{(n)}_{X}$. Hence stopping the EM algorithm at any step does not result in a deterioration of the approximation of the pre-specified feature distribution $Q^{\ast}_{X}$ when compared to the approximation achieved before.
Although the decreasing sequence $\mathrm{KL}_{P_{X}}(h\parallel h_{n})$ converges to a limit $\mathrm{KL}^{\ast}$ for $n\to\infty$, Theorem 3.12 does not exclude the possibility that $\mathrm{KL}^{\ast}>0$ which would suggest that $Q^{\ast}_{X}$ does not correspond to any target distribution related to the source distribution $P$ through label shift. Neither does Theorem 3.12 state that there is a density $h^{\ast}$ with $\mathrm{KL}_{P_{X}}(h\parallel h^{\ast})=\mathrm{KL}^{\ast}$ or that the $Q^{(n)}$ converge to any limit distribution. For the case of a categorical label space, Wu (1983) provided sufficient conditions for the sequence $g_{n}$, $n=0,1,\ldots$, of target prior class probabilities to converge which would entail convergence of the related target joint distributions $Q^{(n)}$.
For the case of a general label space, the following proposition presents consequences of the potential convergence of the target label densities $g_{n}$ which are constructed with the EM algorithm as described by (3.11a) and (3.11b).
Proposition 3.14.
In addition to the assumptions of Theorem 3.12, suppose that for the densities defined in Theorem 3.12 (i) it holds that $\lim_{n\to\infty}g_{n}(y)=g(y)>0$ exists $P_{Y}$-a.s. for $y\in\Omega_{Y}$ and $\lim_{n\to\infty}q^{(n)}_{Y|X=x}(y)=q_{Y|X=x}(y)>0$ exists $P$-a.s. for $(x,y)\in\Omega$. Assume also [^8] The assumptions on the exchangeability of integration and $\lim$ in particular hold true if there is some $M>0$ such that $g_{n}\leq M$ and $q_{Y|X=x}^{(n)}\leq M$ for all $n=0,1,\ldots$ and $x\in\Omega_{X}$. that $\int g(y)\,P_{Y}(dy)=1$,
$\begin{split}\lim\limits_{n\to\infty}\int g_{n}(z)\,P_{Y|X=x}(dz)&=\int g(z)\,P_{Y|X=x}(dz)\quad P_{X}\text{-a.s.\ for}\ x\in\Omega_{X},\quad\text{and}\\ \lim\limits_{n\to\infty}\int q^{(n)}_{Y|X=x}(y)\,h(x)\,P_{X|Y=y}(dx)&=\int q_{Y|X=x}(y)\,h(x)\,P_{X|Y=y}(dx)\quad P_{Y}\text{-a.s.\ for}\ y\in\Omega_{Y}.\end{split}$ (3.14)
Then the following statements hold true:
  • (i)
    $g$ is a density with respect to $P_{Y}$. The joint distribution $\widehat{Q}$ of $X$ and $Y$ defined by $\frac{d\widehat{Q}}{dP}=g(Y)$ is related to $P$ through label shift such that $\frac{d\widehat{Q}_{Y}}{dP_{Y}}=g$ follows. With $\widehat{h}(x)=\frac{d\widehat{Q}_{X}}{dP_{X}}(x)=E_{P}[g(Y)\,|\,X=x]$ for $x\in\Omega_{X}$ it holds that
    $\mathrm{KL}_{P_{X}}(h\parallel\widehat{h})\leq\ \inf_{n}\mathrm{KL}_{P_{X}}(h\parallel h_{n})$ (3.15)
    for the feature densities $h_{n}=E_{P}[g_{n}(Y)\,|\,X=\cdot]$.
  • (ii)
    $g$ is a solution of (3.5c) in the case $\psi=1$. Define $f(x,y)=\frac{g(y)\,h(x)}{\int g(z)\,P_{Y|X=x}(dz)}$ for $(x,y)\in\Omega$ $P$-a.s. Then $f$ is a density with respect to $P$. The joint distribution $Q$ of $X$ and $Y$ defined by $\frac{dQ}{dP}=f$ is related to $P$ through FJS in the sense of Definition 3.1 and it holds that $Q_{X}=Q_{X}^{\ast}$ and $\frac{dQ_{Y}}{dP_{Y}}=g$.
Proposition 3.14 (ii) confirms for general label spaces that if the sequence of densities $g_{n}$ constructed with the EM algorithm converges to a positive limit density $g$, then this density $g$ solves (3.5c) in the case $\psi=1$ and as a consequence also Problem (B) as formulated in Section 2. Thereby, Proposition 3.14 (ii) also confirms for general label spaces the ‘exact fit’ property proven by Tasche (2014, Remark 1) for the case of categorical labels. ‘Exact fit’ as discussed by Tasche (2014) means that there is a target joint distribution such that the ratios of pairs of feature densities conditional on the labels do not depend on feature values (cf. Remark 3.9 above regarding this property for FJS in the case of non-categorical labels) and that the target marginal feature distribution exactly matches a pre-specified or observed feature distribution.
Note that exact fit in the sense of matching a target feature distribution with a constrained FJS (by requiring $\psi=1$ in (3.5c)) is not always possible as pointed out by Titterington et al. (1985, Example 4.3.1) in the context of maximum likelihood estimation of class prior probabilities in the case of binary labels. This observation does not contradict Proposition 3.14 (ii) because in cases of imperfect fit one or more of the conditions listed in the proposition would be violated.
[…]
Does imperfect fit of the pre-specified feature distribution by an approximate solution of Problem (B) matter in practice?
  • The answer is no if it is the target posterior label distribution given the features that is important. This assessment follows from Remark 3.10 that shows that the target posterior feature distribution given the label is the same for label shift and for more general FJS with $\psi=1$ as long as the target label distribution is the same.
  • The answer is yes if it is the target posterior feature distribution given the label that is important. Under label shift (i.e. when the fit is imperfect) the target and source posterior feature distributions given the label are equal. This need not be the case when there is exact fit thanks to FJS with $\psi=1$ as in Proposition 3.14 (ii). This observation is a consequence of Remark 3.9 which demonstrates that the target and source posterior feature distributions given the label may observably differ. See Figure 1 of Tasche (2017) for an example of this phenomenon.
4 Generalized label shift (GLS)
A popular approach to tackling distribution shift (i.e. domain adaptation) is representation learning. The aim with representation learning is to find a mapping of the features into another feature space such that for the joint distribution of the mapped features and the labels the distribution shift is removed or reduced to a more manageable type of shift. Tachet des Combes et al. (2020) proposed with ‘generalized label shift’ (GLS) a variant of representation learning where the representation of the features results in label shift between the source and the target distributions. Wu et al. (2025) use the term ‘Conditionally invariant components’ for GLS and suggest an approach to plausibly identifying such components of the feature variable in the absence of label observations in the target dataset.
He et al. (2021) and Tasche (2022, Proposition 4) noted that GLS implies FJS if the representation mapping is sufficient for the features, i.e. under both the source and the target distribution the label distribution conditional on the feature variable equals the label distribution conditional on the mapped features. The following proposition shows for a sufficient representation mapping that also FJS between the source and the target joint distributions of the mapped features and the labels implies FJS between the original source and target distributions. The function $R$ appearing in Proposition 4.1 stands for the representation mapping. Typically, $R$ would be some function applied to the feature variable $X$, i.e. $R=T(X)$ for some measurable function $T$.
Proposition 4.1.
In the setting of Section 2.1 and under Assumption 2.2, let $R$ be a measurable function on $(\Omega,\mathcal{F})$ such that $\sigma(R)\subset\sigma(X)$ holds. Assume that $P|\sigma(R,Y)$ and $Q|\sigma(R,Y)$ are related through FJS in the sense of Definition 3.1 with $\frac{dQ|\sigma(R,Y)}{dP|\sigma(R,Y)}=\overline{h}(R)\,\overline{g}(Y)$ for measurable non-negative functions $\overline{h}$ and $\overline{g}$. Suppose furthermore that it holds that
$\begin{split}P[Y\in G\,|\,\sigma(X)]&=P[Y\in G\,|\,\sigma(R)],\ \text{and}\\ Q[Y\in G\,|\,\sigma(X)]&=Q[Y\in G\,|\,\sigma(R)]\end{split}$ (4.1)
for all $G\in\mathcal{G}$ (i.e. $\sigma(R)$ is sufficient for $\sigma(X)$ with respect to $\sigma(Y)$ under both $P$ and $Q$ in the sense of Definition 3.1 of Tasche (2021)). In addition, assume that there exists a regular conditional distribution $P_{Y|R}$ of $Y$ with respect to $R$ under $P$ in the sense of Definition 2.1. Then also $P$ and $Q$ are related through FJS with
[…]
Proof of Proposition 4.1.
Observe that $Q[\overline{h}(R)\,\overline{g}(Y)>0]=1=Q\bigl[\overline{h}(R)\,E_{P}[\overline{g}(Y)\,|\,\sigma(R)]>0\bigr]$ (proof like in Lemma A1 of Tasche, 2022). This implies
[…]
Thanks to the existence of $P_{Y|R}$, (4.1) implies $E_{P}[T(Y)\,|\,\sigma(R)]=E_{P}[T(Y)\,|\,\sigma(X)]$ for all $\mathcal{G}$-Borel measurable functions $T\geq 0$. Therefore, it follows from (4.3a) that
$0\ =\ Q\bigl[E_{P}[\overline{g}(Y)\,|\,\sigma(X)]=0\bigr]\ =\ E_{P}\bigl[E_{P}[f(X,Y)\,|\,\sigma(X)]\,\mathbf{1}_{\{E_{P}[\overline{g}(Y)\,|\,\sigma(X)]=0\}}\bigr].$ (4.3b)
(4.3a) and (4.3b) together with the sufficiency assumptions and the generalized Bayes formula imply that the following calculation is valid for all $H\in\mathcal{H}$ and $G\in\mathcal{G}$:
$Q[X\in H,Y\in G]$ $=Q\bigl[\mathbf{1}_{\{X\in H\}}\,Q[Y\in G\,|\,\sigma(X)]\bigr]$
$=Q\bigl[\mathbf{1}_{\{X\in H\}}\,Q[Y\in G\,|\,\sigma(R)]\bigr]$
$=E_{Q}\left[\mathbf{1}_{\{X\in H\}}\,\frac{E_{P}[\overline{g}(Y)\,\mathbf{1}_{\{Y\in G\}}\,|\,\sigma(R)]}{E_{P}[\overline{g}(Y)\,|\,\sigma(R)]}\right]$
$=E_{P}\left[f(X,Y)\,\mathbf{1}_{\{X\in H\}}\,\frac{E_{P}[\overline{g}(Y)\,\mathbf{1}_{\{Y\in G\}}\,|\,\sigma(R)]}{E_{P}[\overline{g}(Y)\,|\,\sigma(R)]}\right]$
$=E_{P}\left[E_{P}[f(X,Y)\,|\,\sigma(X)]\,\mathbf{1}_{\{X\in H\}}\,\frac{E_{P}[\overline{g}(Y)\,\mathbf{1}_{\{Y\in G\}}\,|\,\sigma(R)]}{E_{P}[\overline{g}(Y)\,|\,\sigma(R)]}\right]$
$=E_{P}\left[\mathbf{1}_{\{X\in H\}\cap\{Y\in G\}}\,\frac{E_{P}[f(X,Y)\,|\,\sigma(X)]}{E_{P}[\overline{g}(Y)\,|\,\sigma(X)]}\,\overline{g}(Y)\right].$
This implies (4.2). ∎
Recall that if $P|\sigma(R,Y)$ and $Q|\sigma(R,Y)$ are related through label shift, then according to Definition 3.1 of Tachet des Combes et al. (2020) $P$ and $Q$ are related through generalized label shift (GLS).
Corollary 4.2.
Under the assumptions of Proposition 4.1, suppose that $P|\sigma(R,Y)$ and $Q|\sigma(R,Y)$ are related through label shift in the sense of Definition 3.2 (ii). Denote by $g=\frac{dQ_{Y}}{dP_{Y}}$ the label density of $Q_{Y}$ with respect to $P_{Y}$. Then $P$ and $Q$ are related through FJS with
[…]
Proof of Corollary 4.2..
Corollary 4.2 is a special case of Proposition 4.1 since Proposition 3.3 and its proof imply that $\frac{dQ|\sigma(R,Y)}{dP|\sigma(R,Y)}=g(Y)$. Replacing $\overline{g}$ in (4.2) with $g$ yields (4.4). ∎
Corollary 4.2 suggests an alternative to domain adaptation through representation learning in combination with a GLS assumption. One of the objectives of representation learning is avoiding losing predictive power for classification or regression in comparison to classification or regression based on the original features. For this reason, the sufficiency condition (4.1) of Proposition 4.1 is likely to be satisfied at least approximately.
But then by Corollary 4.2, if $Q$ and $P$ are related through GLS they also are related through FJS, even with the same density of the features conditional on the label for all label values as discussed in Remark 3.9, i.e. $\psi=1$ in (3.5c). This observation applies independently of the representation mapping $R$. Hence it might make sense to renounce finding a representation $R$ for achieving GLS and instead resort to the EM algorithm for solving Problem (B) as presented in Section 3.2.3.
5 Concluding remarks
In this paper, we have extended the notion of factorizable joint shift (FJS) to cover the regression case with general label space and shed more light on the relation between FJS, covariate shift and label shift. We also have revisited the property of generalized label shift (GLS) to imply FJS if the involved representation mapping is sufficient for the original features. We have shown that this property is preserved in settings with general label spaces.
In order to facilitate the related analyses, we have introduced an assumption on the source distribution with regard to the feasibility of regular conditional distributions of the labels given the features and vice versa. Another assumption on the existence of a density of the source distribution with respect to the product measure of its feature and label marginal distributions has turned out to be useful for the reconciliation of the theoretical results with results for categorical labels as well as for calculations in practice.
With regard to fitting FJS to given target feature and label marginal distributions, we have presented an approach based on a non-linear integral equation for which we have proposed an iterative solution. Regarding fitting FJS to a given target feature distribution without recourse to any label observations, in order to cover the case of a general label space we have presented a generalisation of the expectation maximisation (EM) approach for estimating target prior class probabilities (Saerens et al., 2002) and shown that it is fit for purpose.
Future research topics could be the computational properties of the estimation approaches proposed in Sections 3.2.1 and 3.2.3 and the study of other specific types of distribution shift (e.g. sparse joint shift, Chen et al., 2022) in the face of general label spaces.
Appendix A Appendix: Density estimation as binary classification problem
It is a well-known fact that a Radon-Nikodym density of a probability measure with respect to another probability measure can be represented by means of the posterior probability in a binary classification problem (Qin, 1998; Sugiyama et al., 2012, Chapter 4). With the following proposition, we provide a formal statement and proof of this fact.
Proposition A.1.
Let $(\overline{M},\mathcal{M})$ be a measurable space and $\mathbb{P}_{0}$ and $\mathbb{P}_{1}$ be probability measures on $(\overline{M},\mathcal{M})$. Assume that $\lambda\geq 0$ is a density of $\mathbb{P}_{1}$ with respect to $\mathbb{P}_{0}$, i.e. it holds that $\lambda=\frac{d\mathbb{P}_{1}}{d\mathbb{P}_{0}}$.
Define $\overline{\Omega}=\{0,1\}\times\overline{M}\times\overline{M}$ and the $\sigma$-algebra $\overline{\mathcal{F}}$ on $\overline{\Omega}$ by $\overline{\mathcal{F}}=\mathfrak{P}(\{0,1\})\otimes\mathcal{M}\otimes\mathcal{M}$ (the product $\sigma$-algebra of $\mathfrak{P}(\{0,1\})$, $\mathcal{M}$ and $\mathcal{M}$).
For $\omega=(s,m_{0},m_{1})\in\overline{\Omega}$ denote by $S$, $X_{0}$ and $X_{1}$ the coordinate projections, i.e. $S(\omega)=s$, $X_{0}(\omega)=m_{0}$ and $X_{1}(\omega)=m_{1}$.
Let $0<p<1$ be fixed and define the probability measure $\mathbb{P}_{p}$ on $(\overline{\Omega},\overline{\mathcal{F}})$ as the product measure such that $S$, $X_{0}$ and $X_{1}$ are independent under $\mathbb{P}_{p}$ with $\mathbb{P}_{p}[S=1]=p=1-\mathbb{P}_{p}[S=0]$, $\mathbb{P}_{p}[X_{0}\in M]=\mathbb{P}_{0}[M]$ and $\mathbb{P}_{p}[X_{1}\in M]=\mathbb{P}_{1}[M]$ for $M\in\mathcal{M}$.
In addition, define $X:\overline{\Omega}\to\overline{M}$ by $X=\begin{cases}X_{0},&\text{if}\ S=0\\ X_{1},&\text{if}\ S=1\end{cases}$ and $\eta_{p}(m)=\mathbb{P}_{p}[S=1\,|\,X=m]$.
Then it holds that
$\eta_{p}(m)\ =\ \frac{p\,\lambda(m)}{p\,\lambda(m)+1-p},\quad\mathbb{P}_{0}\text{-a.s.}\ m\in\overline{M}.$ (A.1)
Proof of Proposition A.1.
On the one hand, by the definition of conditional probability, it follows for $M\in\mathcal{M}$ that
$\mathbb{P}_{p}[S=1,X\in M]$ $=E_{\mathbb{P}_{p}}\bigl[\eta_{p}(X)\,\mathbf{1}_{\{X\in M\}}\bigr]$
[…]
$=(1-p)\,E_{\mathbb{P}_{0}}[\eta_{p}\,\mathbf{1}_{M}]+p\,E_{\mathbb{P}_{1}}[\eta_{p}\,\mathbf{1}_{M}]$
$=E_{\mathbb{P}_{0}}\bigl[\mathbf{1}_{M}\,\eta_{p}\,(p\,\lambda+1-p)\bigr].$ (A.2)
On the other hand, the definition of $\mathbb{P}_{p}$ implies
$\mathbb{P}_{p}[S=1,X\in M]\ =\ \mathbb{P}_{p}[S=1,X_{1}\in M]\ =\ p\,\mathbb{P}_{1}[M]\ =\ p\,E_{\mathbb{P}_{0}}[\lambda\,\mathbf{1}_{M}].$
Together with (A.2), this observation implies $\eta_{p}(m)\,(p\,\lambda(m)+1-p)=p\,\lambda(m)$ $\mathbb{P}_{0}$-almost surely for $m\in\overline{M}$. Since $p\,\lambda(m)+1-p\geq 1-p>0$, (A.1) follows. ∎
Some consequences of Proposition A.1 are as follows:
  • (A.1) implies $\eta_{p}(m)<1$ for $m\in\overline{M}$ $\mathbb{P}_{0}$-a.s. and therefore also
    $\lambda(m)\ =\ \frac{1-p}{p}\,\frac{\eta_{p}(m)}{1-\eta_{p}(m)}.$ (A.3)
  • (A.3), in particular, implies that the parameter $p$ is eliminated when (A.1) is solved for $\lambda(m)$.
  • Suppose that there are two separate samples of realisations of a variable (with an arbitrary but fixed range of values), and that the two samples have been generated under different distributions of the variable in question. Assume further that one of these distributions is absolutely continuous with respect to the other distribution such that there exists a density of the former distribution with respect to the latter distribution. Proposition A.1 can be used to estimate the density in the following way:
    1. 1)
      Combine the two samples and assign the label 0 to the sample elements which come from the reference distribution for the density and the label $1$ to the sample elements which stem from the distribution which is absolutely continuous with respect to the other density.
    2. 2)
      Train a binary probabilistic binary classifier $\eta$ to classify the elements of the combined sample according to their labels.
    3. 3)
      Let $p=\frac{\text{\# labels 1}}{\text{size of combined sample}}$.
    4. 4)
      Use (A.3) to represent the density in terms of the learnt probabilistic classifier $\eta$ and $p$ from step 3).
(A.3) shows that over-sampling and under-sampling techniques may be deployed when training the classifier without changing the resulting density.
Consider the case where the density $\varphi$ of $P=P_{(X,Y)}$ with respect to $P_{X}\otimes P_{Y}$ according to Assumption 2.9 is estimated by means of binary classification as described above. Then one sample from $P$ suffices because a sample for $P_{X}\otimes P_{Y}$ can be generated as the set of pairs of each element of the marginal sample for $X$ and each element of the marginal sample for $Y$. But this method creates strong imbalance of the labels in the classification sample. Therefore, the robustness of the classification approach with respect to under- and over-sampling becomes important.
Appendix B Appendix: Proofs for Section 3.2.3
Proof of Theorem 3.12..
The proof of (i) works by means of induction over $n$. The assertions for $q^{(n)}_{Y|X=x}$ immediately follow from the assertions for $g_{n}$ because $g_{n}$ in (3.11a) is normalised to become a density. Assume that the assertions for $g_{n}$ and $q^{(n)}_{Y|X=x}$ are true and consider $g_{n+1}$. By (3.11b), $g_{n+1}>0$ follows from $q^{(n)}_{Y|X=x}>0$ and $h>0$. The density property $\int g_{n+1}\,dP_{Y}=1$ is the result of the following calculation which makes use of Fubini’s theorem:
$\int g_{n+1}\,dP_{Y}$ $=\int\int q_{Y|X=x}^{(n)}(y)\,h(x)\,P_{X|Y=y}(dx)\,P_{Y}(dy)$
$=\int\int q_{Y|X=x}^{(n)}(y)\,h(x)\,P_{Y|X=x}(dy)\,P_{X}(dx)$
$=\int\frac{h(x)}{\int g_{n}(z)\,P_{Y|X=x}(dz)}\int g_{n}(y)\,P_{Y|X=x}(dy)\,P_{X}(dx)$
$=\int h\,dP_{X}=1.$
Regarding (ii), for any fixed non-negative integer $k$, by (3.11a) and (3.4a) it holds that $q^{(k)}_{Y|X=x}=\frac{dQ^{(k)}_{Y|X=x}}{dP_{Y|X=x}}$ as the choice of notation already suggested. By definition of the conditional density (as in Theorem 2.6) and of $Q^{(k)}$, it follows that $\log\bigl(h_{k}(x)\bigr)=\log\bigl(g_{k}(y)\bigr)-\log\bigl(q^{(k)}_{Y|X=x}(y)\bigr)$ for $(x,y)\in\Omega$. Taking the expectation with respect to $Q^{(n)}_{Y|X=x}$ on both sides of the equation for a possibly different non-negative integer $n$ yields for $x\in\Omega_{X}$
$\log\bigl(h_{k}(x)\bigr)$ $=\int\log\bigl(g_{k}(y)\bigr)\,Q^{(n)}_{Y|X=x}(dy)-\int\log\bigl(q^{(k)}_{Y|X=x}(y)\bigr)\,Q^{(n)}_{Y|X=x}(dy)$
$=\int\log\bigl(g_{k}(y)\bigr)\,q^{(n)}_{Y|X=x}(y)\,P_{Y|X=x}(dy)-\int\log\bigl(q^{(k)}_{Y|X=x}(y)\bigr)\,q^{(n)}_{Y|X=x}(y)\,P_{Y|X=x}(dy),$
again by the definition of conditional density. Now multiply both sides of the last equation with $h(x)$ and take the expectation with respect to $P_{X}$, then apply Fubini’s theorem and make use of (3.11b) to obtain
$\int h(x)\,\log\bigl(h_{k}(x)\bigr)P_{X}(dx)$ $=\int\log\bigl(g_{k}(y)\bigr)\left(\int q^{(n)}_{Y|X=x}(y)\,h(x)\,P_{X|Y=y}(dx)\right)P_{Y}(dy)$
$\qquad-\int h(x)\left(\int\log\bigl(q^{(k)}_{Y|X=x}(y)\bigr)\,q^{(n)}_{Y|X=x}(y)\,P_{Y|X=x}(dy)\right)P_{X}(dx)$
$=\int\log\bigl(g_{k}(y)\bigr)g_{n+1}(y)P_{Y}(dy)$ (B.1)
$\qquad-\int h(x)\left(\int\log\bigl(q^{(k)}_{Y|X=x}(y)\bigr)\,q^{(n)}_{Y|X=x}(y)\,P_{Y|X=x}(dy)\right)P_{X}(dx).$
From Jensen’s inequality follows that
$\int\log\bigl(g_{n+1}(y)\bigr)g_{n+1}(y)P_{Y}(dy)$ $\geq\int\log\bigl(g_{n}(y)\bigr)g_{n+1}(y)P_{Y}(dy)\quad\text{and}$
$\int\log\bigl(q^{(n+1)}_{Y|X=x}(y)\bigr)\,q^{(n)}_{Y|X=x}(y)\,P_{Y|X=x}(dy)$ $\leq\int\log\bigl(q^{(n)}_{Y|X=x}(y)\bigr)\,q^{(n)}_{Y|X=x}(y)\,P_{Y|X=x}(dy),\ \text{for}\ x\in\Omega_{X}.$
These inequalities together with (B.1) both for $k=n+1$ and $k=n$ imply $\int h(x)\,\log\bigl(h_{n+1}(x)\bigr)P_{X}(dx)\geq\int h(x)\,\log\bigl(h_{n}(x)\bigr)P_{X}(dx)$ and by subtracting $\int h(x)\,\log\bigl(h(x)\bigr)P_{X}(dx)$ on both sides of this inequality also (3.13). ∎
Proof of Proposition 3.14..
The assumption $\int g\,dP_{Y}=1$ implies that $g$ is a density with respect to $P_{Y}$. Hence $\widehat{Q}$ is well-defined and by Proposition 3.3 $\widehat{Q}$ and $P$ are related through label shift, with the consequence that $g$ is the density of $\widehat{Q}_{Y}$ with respect to $P_{Y}$ and $\widehat{h}$ the density of $\widehat{Q}_{X}$ with respect to $P_{X}$. Since $\int g\,dP_{Y}=1=\int g_{n}\,dP_{Y}$ for all $n$ the sequence $g_{n}(Y)$, $n=0,1,\ldots$, is uniformly integrable under $P$ and consequently $g_{n}(Y)\to g(Y)$ $P$-a.s. implies $\lim_{n\to\infty}E_{P}\bigl[|g_{n}(Y)-g(Y)\bigr]=0$. These observations imply that also $h_{n}(X)=E_{P}[g_{n}(Y)\,|\,\sigma(X)]$, $n=0,1,\ldots$, is uniformly integrable and that $\lim_{n\to\infty}E_{P}\bigl[|h_{n}(X)-\widehat{h}(X)\bigr]=0$. However, from uniform integrability and L1-convergence follows $\lim_{n\to\infty}h_{n}(x)=\widehat{h}(x)$ $P_{X}$-a.s. By a combination of (3.13) from Theorem 3.12 (ii) and Fatou’s lemma the almost sure convergence of $h_{n}$ implies (3.15).
Regarding Proposition 3.14 (ii), observe that by $g>0$, (3.11a), (3.11b) and (3.14) it follows for $(x,y)\in\Omega$ $P$-a.s. that
$g(y)$ $=\int\frac{g(y)}{\int g(z)\,P_{Y|X=x}(dz)}\,h(x)\,P_{X|Y=y}(dx)$
$\iff\ 1$ $=\int\frac{h(x)}{\int g(z)\,P_{Y|X=x}(dz)}\,P_{X|Y=y}(dx).$ (B.2)
By (B.2) $g$ solves (3.5c) in the case $\psi=1$. Define $\overline{h}(x)=\frac{h(x)}{\int g(z)\,P_{Y|X=x}(dz)}$ and $\overline{g}(y)=g(y)$. Then taking into account (2.4b), (B.2) and Theorem 3.5 (ii) imply the assertion. ∎